### v1

# MoGe: Unlocking Accurate Monocular Geometry Estimation for Open-Domain Images with Optimal Training Supervision

Abstract
We present MoGe, a powerful model for recovering 3D geometry from monocular open-domain images. Given a single image, our model directly predicts a 3D point map of the captured scene with an affine-invariant representation, which is agnostic to true global scale and shift. This new representation precludes ambiguous supervision in training and facilitate effective geometry learning. Furthermore, we propose a set of novel global and local geometry supervisions that empower the model to learn high-quality geometry. These include a robust, optimal, and efficient point cloud alignment solver for accurate global shape learning, and a multi-scale local geometry loss promoting precise local geometry supervision. We train our model on a large, mixed dataset and demonstrate its strong generalizability and high accuracy. In our comprehensive evaluation on diverse unseen datasets, our model significantly outperforms state-of-the-art methods across all tasks, including monocular estimation of 3D point map, depth map, and camera field of view. Code and models will be released on our project page.
{strip}[Uncaptioned image]
Figure 1: Given any image, our method reconstructs an affine-invariant 3D point map of the scene and can also produce a depth map and the camera focal length. The model yields high-quality shapes and generalizes well to open-domain images. (Best view with zoom)
1 Introduction
Estimating the 3D geometry of general scenes is a fundamental task in computer vision. While 3D reconstruction from a set of images has been extensively studied with Structure-from-Motion (SfM) [1, 47] and Multi-View Stereo (MVS) [48, 64] techniques, recovering the 3D geometry from a single image in an arbitrary domain remains a significant challenge due to its highly ill-posed nature.
For monocular geometry estimation (MGE), a common approach involves first estimating a depth map, subject to an unknown scale (and shift), and then combining it with camera intrinsics to recover the 3D shape via unprojection [69, 41]. For the former, significant advancements in monocular depth estimation (MDE) have been achieved in recent years by training on large-scale data [71, 65, 29, 20]. However, inferring camera intrinsics (e.g., focal length) from single images remains challenging due to the high degree of ambiguity when strong geometric cues are absent. Inaccurate camera parameters can lead to significant geometry distortions even when used with ground truth depth maps. Recently, DUSt3R [57] proposed predicting a 3D point map from an image by mapping each pixel to a free 3D point, thereby bypassing camera parameter estimation during geometry recovery. Although the model can estimate monocular geometry, it is primarily designed for multi-view scenarios and is trained on image pair inputs.
In this paper, we introduce a new direct 3D geometry estimation method designed for single images in the open domain. Our model architecture is simple and straightforward: it directly predicts point maps from images, which can further derive depth map and camera focal length or FOV if needed. Different from DUSt3R [57] which uses scale-invariant point maps, our method predicts affine-invariant point maps, where the 3D points are subject to an unknown global scale as well as a 3D shift. This alteration is important as it eliminates the focal-distance ambiguity which is detrimental to the network training – we provide intuitive explanations with illustration as well as empirical evidence to support this advancement.
Figure 2: Accuracy ranking of existing methods and ours in our extensive evaluation on monocular 3D point map (scale-invariant and affine-invariant), depth map (scale-invariant depth, affine-invariant depth, and affine-invariant disparity), and camera field of view. Outer methods rank higher. See our experiments for details.
More importantly, our findings indicate that the design of training supervisions is crucial to achieve robust and accurate point map prediction from single images. Similar to previous MDE approaches, we need to compute a global scaling factor and a translation to align the predicted results with the ground truth during training. However, previous methods for this global alignment calculation are either sensitive to outliers or solved with coarse approximations, leading to unsatisfactory supervision. We propose a Robust, Optimal, and Efficient (ROE) global alignment solver to resolve scale and shift for the affine-invariant pointmap loss, which substantially improved the training effectiveness and final accuracy. On the other hand, the effective learning of local, region-wise geometry has been largely neglected previously. In monocular geometry estimation, the relative distance between different objects can be ambiguous, which hinders precise local geometry learning when a global alignment computed with all objects is applied. In light of this, we propose a multi-scale local geometry loss, which imposes penalties on the local discrepancies of 3D point clouds, each under independent and optimal affine alignment. This design significantly boosted the accuracy of our local geometry prediction.
We train our model on a large-scale training corpus sourced from various existing datasets. Our model demonstrates strong generalizability and accuracy in monocular geometry estimation for open-domain images (some examples are shown in Fig. 1). The zero-shot evaluation on eight unseen datasets shows that it outperforms previous MGE methods that output point clouds by a large margin (over 35% error reduction compared to the previous best). Moreover, our method also significantly outperforms previous approaches that focus on the sub-tasks of MDE and camera FoV (20$\sim$30error reduction for the former and over 20% for the latter). Figure 2 shows the performance ranking of different methods in our evaluation, where our method ranks top across all tasks/metrics and among all compared methods.
Our contributions are summarized as follows: 1) We propose a new direct MGE method for open-domain images using affine-invariant point maps. 2) We establish a set of novel and effective global and local supervisions for robust and precise geometry recovery. 3) We demonstrate the substantial enhancement in performance of our model over existing methods on MGE, MDE, and camera FOV estimation across diverse datasets. We hope that our method can serve as a strong and versatile foundational model for solving monocular geometry problems, facilitating applications such as 3D-aware image editing, image relighting, depth-to-image synthesis, novel view synthesis, and 3D scene understanding, as well as providing initial geometrical priors for video-based or multiview-based 3D reconstruction.
2 Related Work
Monocular depth estimation.
As a crucial precursor to monocular geometry estimation, monocular depth estimation is a long-standing task and has been extensively studied in the past. A number of methods [18, 4, 71, 26, 41, 5, 35] have been developed to estimate depth with metric scale, which often rely heavily on data from specific sensors such as RGBD cameras, LiDAR, and calibrated stereo cameras. This dependence restricts their applicability to specific domains, such as indoor and driving scenes. In contrast, relative depth estimation has garnered significant attention due to its ability to utilize a much wider variety of data, thereby exhibiting enhanced generalizability. A prevalent approach is to predict relative depth in an affine-invariant manner, either through direct regression [10, 34, 22, 11, 68, 42, 65, 66] or generative modeling [29, 20, 54, 23]. Despite the recent advancements in monocular depth estimation, recovering 3D shape from depth information always necessitates known camera intrinsics.
Monocular point map estimation.
Compared to depth map, 3D point map offers a more straightforward representation of 3D geometry. Monocular point map estimation aims to recover free 3D points for each pixel on a single image. Several approaches tackle this challenge by predicting camera parameters alongside depth estimation. For example, LeReS [69, 70] introduced a two-stage pipeline with an affine-invariant depth prediction module followed by a point cloud module for recovering shift and camera focal length. UniDepth [41] proposed a self-promptable camera module that predicts dense camera representation to condition the subsequent depth estimation module. DUSt3R [57] employs a more straightforward end-to-end model to directly map two-view images to camera space point maps, and it can adapt to monocular scenarios by using two identical images as input. It uses the camera-space scale-invariant pointmap which incorporates camera shift information. This representation can suffer from the focal-distance ambiguity and jeopardize pointmap learning. In contrast, our approach is designed to handle monocular input and we employ affine-invariant point map with meticulously designed training supervisions for more effective geometry learning.
Camera intrinsics estimation.
Estimating camera intrinsics is a fundamental task in 3D vision. While traditional multiview camera calibration using special patterns is well studied [73], estimating camera intrinsics from single images is ill-posed and remains challenging. For the latter, earlier works utilized known 3D shapes [59] or vanishing points [13] to solve the problem. More recently, learning-based approaches have been proposed [61, 7, 32, 28, 75] to handle in-the-wild images, but the results are far from satisfactory. In this work, we demonstrate that our camera parameter estimation derived from point maps can achieve state-of-the-art accuracy and offer remarkable generalizability.
Large-scale data training for monocular geometry.
Recently, there has been an emerging trend [42, 17, 5, 41, 71, 26, 57, 65, 66] in monocular geometry estimation and depth estimation that leverages large-scale datasets coupled with advanced network backbones [15, 43, 40] to enhance performance. MiDaS [42] marks an important milestone by mixing large datasets from various domains for training. Depth Anything [65] leverages both labeled and extensive unlabeled data to improve generalization. Its successive work [66] further highlights the importance of high-quality synthetic data for capturing rich shape details in depth prediction. These advancements underscore the vital importance of large-scale training data in improving the performance of monocular geometry estimation.
Figure 3: Method overview. Our model consists of a ViT encoder and a convolutional decoder. It predicts an affine-invariant point map as well as a mask that excludes regions with undefined geometry (e.g., infinity). Depth, camera shift, and focal length can be further derived from the model output. For training, we design robust and effective supervisions focusing on both the global and local geometry.
Figure 4: The focal-distance ambiguity and effects of different 3D point representations. (a) For similar images captured with varying camera focal length and distance to the objects, perceiving their true camera setup is challenging and models often produce similar geometries. (b) Inconsistent supervision signals occur with only scale alignment. (c) Consistent geometry supervision with an additional translation alignment.
Figure 5: The relative distance ambiguity, where distances between objects can be hard to predict.
3 Approach
Our method takes a single image as input and directly predicts the geometry of the scene represented by 3D points. An overview of the method is shown in Fig. 3.
3.1 Affine-invariant point map
For an image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, our model $F_{\theta}$ infers the 3D coordinates of the image pixels, represented by a point map $\mathbf{P}\in\mathbb{R}^{H\times W\times 3}$, i.e.,:
$F_{\theta}:\mathbf{I}\mapsto\mathbf{P}.$ (1)
The $X$ and $Y$ axes of the coordinate frame of $\mathbf{P}$ are aligned with the $u$ and $v$ axes in the image space respectively.
Monocular geometry estimation often suffers from focal-distance ambiguity, as shown in Fig. 5 (a). To deal with this issue, we propose to predict affine-invariant point map $\mathbf{P}$, i.e., $\mathbf{P}$ is agnostic to the global scale $s\in\mathbb{R}$ and offset $\mathbf{t}\in\mathbb{R}^{3}$ and thus $\mathbf{P}\cong s\mathbf{P}+\mathbf{t},\forall s,\forall\mathbf{t}$. Compared to the scale-invariant supervision [57] shown in Fig. 5 (b), the discrepancies in focal lengths and camera positions between the predictions and ground truth are amended by a translation, as shown in Fig. 5 (c). This ensures consistent geometry supervision when focal length is ambiguous, enabling more effective geometry learning.
Following accepted practice, we assume that the camera principal point coincides with image center and pixels are square. In this case, $\mathbf{t}$ can be simplified with a $Z$-axis shift $t_{z}$ (i.e., $t_{x}=t_{y}=0$).
Recovering camera focal and shift.
The affine-invariant point map representation can be used to recover both camera shift and focal length. Given the predicted 3D points $(x_{i},y_{i},z_{i})$ and their corresponding 2D pixels $(u_{i},v_{i})$, we solve for camera focal length $f$ and $Z$-axis shift by minimizing the projection error:
$\min_{f,t_{z}^{\prime}}\sum_{i=1}^{N}\left({fx_{i}\over z_{i}+t_{z}^{\prime}}-u_{i}\right)^{2}+\left({fy_{i}\over z_{i}+t_{z}^{\prime}}-v_{i}\right)^{2},$ (2)
where $t_{z}^{\prime}=t_{z}/s$. Eq. 2 can be efficiently solved with a fast iterative optimization algorithm, which converges within 10 iterations, taking approximately $3$ms. More details about the solution can be found in the suppl. materials.
With $t_{z}^{\prime}$ recovered, the scale-invariant depth map and point map in the camera space can be obtained by adding it to the $z$ coordinates.
3.2 Training Objectives
Global point map supervision.
Let $\hat{\mathbf{p}}_{i}$ denote the predicted 3D point for the $i$-th pixel, and $\mathbf{p}_{i}$ the corresponding ground truth. The global point map loss is defined as:
$\mathcal{L}_{G}=\sum_{i\in{\mathcal{M}}}{\frac{1}{{z}_{i}}\|s\hat{\mathbf{p}}_{i}+\mathbf{t}^{*}-{\mathbf{p}}_{i}\|_{1}},$ (3)
where $\|\!\cdot\!\|_{1}$ denotes the $L_{1}$ norm, $s$ and $\mathbf{t}$ are the alignment parameters that transform the predicted affine-invariant point map to the ground-truth camera space, and ${\mathcal{M}}$ is the mask for regions with labels. The weighting term $\frac{1}{{z}_{i}}$, where $z_{i}$ is the $z$-coordinate of ${\mathbf{p}}_{i}$, is applied to balance the supervision signal across extreme depths variations.
To apply the global loss $\mathcal{L}_{G}$ for training, we need to first determine $s$ and $\mathbf{t}$. Previous methods for affine-invariant depth estimation often use coarse approximations for this alignment. For example, a predominate strategy is to use the median depth values of the two sets as the anchor to compute shift followed by scaling factor calculation [42, 65, 66]. The range normalization strategy used by [29, 20] can be viewed as computing shift using minimal or a predefined fraction of depth values. All these simple strategies are non-optimal for the problem setup and could lead to unsatisfactory supervision.
In this work, we propose an optimal solver to the alignment parameters. Specifically, we determine $s^{*}$ and $\mathbf{t}^{*}$ via
$\begin{split}(s^{*},\mathbf{t}^{*})=\mathop{\arg\!\min}_{s,\mathbf{t}}\sum_{i\in{\mathcal{M}}}\frac{1}{{z}_{i}}\left\|s\hat{\mathbf{p}}_{i}+\mathbf{t}-{\mathbf{p}}_{i}\right\|_{1},\end{split}$ (4)
where $t_{x}$ and $t_{y}$ are 00. One possible approach to solve this equation is to formulate it as an absolute residuals optimization problem and solve it using linear programming techniques such as simplex algorithms and the interior-point method [8]. However, these methods are known to have high computational complexity, often exceeding $O(N^{3})$. Given thousands of points, a typical implementation may require computation time of a few seconds or even higher, which is inefficient for network training.
Instead, we develop an efficient parallelized searching algorithm to solve it optimally. Our theoretical analysis (Sec. A.2) demonstrates that the minimum must occur when $s\hat{z}^{(k)}+t_{z}-z^{(k)}=0$ for some $k=1,\ldots,N,\ k\in\mathcal{M}$. This insight allows us to substitute $t_{z}$ with $s$ and decompose the problem into a series of one-dimensional subproblems with scale variable only, which can be solved in parallel. The overall computational complexity is $O(N^{2}\log N)$ and efficient training is achieved on GPU. A high-level description of this solution is presented in Algorithm 1 and more details can be found in the suppl. materials.
In practice, we noticed that Eq. 4 is still sensitive to outliers even if the $L_{1}$ error norm has been used. If the model wrongly predicts a close foreground edge pixel as a far background point, the objective will be dominated by the incorrect edge prediction. To further improve the robustness, we apply a clipping operator $\operatorname{clip}\left(\cdot,\tau\right)=\max\big{(}\min(\cdot,\tau),-\tau\big{)}$ on the per-point residuals, which brings accuracy gains as we will show in the experiments. Adding the clipping operator makes the problem non-convex, but our optimal solver is still applicable with minor modifications.
In summary, our alignment solver is made Robust, Optimal, and Efficient, and hence we call it the ROE solver.
Algorithm 1 Optimal alignment solver to Eq. 4
function solve_subproblem
    Enumerating non-differentiable points of objective function,
         at zero per-coordinate alignment residuals.
    Determining extrema by derivative computation.
    Return $s^{(k)}$ with smallest objective value $l^{(k)}$ at extrema.
end function
for index $k=1,...,N$ do $\triangleright$ parallel computation
    Formulate subproblem by substituting $t_{z}$ with $z^{(k)}\!-\!s\hat{z}^{(k)}$.
    Solve scale $s^{(k)}$ and $l^{(k)}$ via SOLVE_SUBPROBLEM.
    Obtain translation $t_{z}^{(k)}$ as $z^{(k)}\!-\!s^{(k)}\hat{z}^{(k)}$.
end for
Select optimal $s^{*}$ and $t_{z}^{*}$ with smallest function value $l^{(k)}$.
Multi-scale local geometry loss.
In monocular geometry estimation, the relative distance between different objects can be ambiguous and difficult to predict, as illustrated in Fig. 5. This hinders precise local geometry learning when a global alignment computed with all objects is applied. To enhance the supervision for local geometry, we propose a loss function which measures the accuracy of local sphere regions with independent alignment.
Specifically, given a ground-truth 3D point ${\mathbf{p}}_{j}$ as the anchor, we first select the point set within the spherical region centered at ${\mathbf{p}}_{j}$, defined as:
${\mathcal{S}}_{j}=\{i\mid\|{\mathbf{p}}_{i}-{\mathbf{p}}_{j}\|\leq r_{j},i\in{\mathcal{M}}\},$ (5)
where $r_{j}$ is the radius. We set $r_{j}=\alpha\cdot z_{j}\cdot\frac{\sqrt{W^{2}+H^{2}}}{2\cdot f}$ where $z_{j}$ is the $z$-coordinate of $p_{j}$, $f$ is the focal length, and $W$ and $H$ are image width and height. This way, the hyper-parameter $\alpha\in(0,1)$ approximates the proportion of the projected sphere’s diameter to the image diagonal.
Then, we apply the aforementioned ROE alignment solver to align the point maps and compute error. For each sphere scale parameter $\alpha$, we sample a set of anchor points $\mathcal{H}_{\alpha}$ and compute the loss as
$\mathcal{L}_{S(\alpha)}=\sum_{j\in\mathcal{H}_{\alpha}}l_{{\mathcal{S}}_{j}}=\sum_{j\in\mathcal{H}_{\alpha}}\sum_{i\in{S}_{i}}\frac{1}{z_{i}}\left\|s^{*}\hat{\mathbf{p}}_{i}+\mathbf{t}^{*}-{\mathbf{p}}_{i}\right\|_{1}$ (6)
In practice, we use three $\alpha$ scales ranging from coarse to fine: $1/4$, $1/16$, and $1/64$, denoted as $\mathcal{L}_{S_{1}}$, $\mathcal{L}_{S_{2}}$ and $\mathcal{L}_{S_{3}}$, respectively.
Normal loss.
For better surface quality, we additionally supervise the normal computed from the predicted point map with the ground truth:
$\mathcal{L}_{N}=\sum_{i\in{\mathcal{M}}}\angle(\hat{\mathbf{n}}_{i},\mathbf{n}_{i})$ (7)
where the normal $\mathbf{n}_{i}$ of pixel $i$ is obtained from the cross product of its adjacent edges on the image grid.
Mask loss.
The distance to infinity region (e.g., sky) in outdoor scenes and plain backgrounds in object-only images are undefined. We add a single-channel head to our model for predicting the mask $\hat{\mathbf{M}}\in\mathbb{R}^{H\times W}$ where the distance is valid:
$\mathcal{L}_{M}=\|\hat{\mathbf{M}}-(1-{\mathbf{M}}_{\text{inf}})\|_{2}^{2}$ (8)
where ${\mathbf{M}}_{\text{inf}}$ is the infinity mask label. For synthetic data, the ground truth masks are readily available, while for real outdoor scenes, we use SegFormer [62] to obtain sky masks.
At inference time, we use a threshold of $0.5$ to binarize the mask prediction. Recovery camera focal and shift involves masked points only, excluding the regions of meaningless geometry.
3.3 Training data
Recently, numerous synthetic and real RGBD datasets have been utilized in large-scale training for monocular geometry estimation. In this work, we collected $21$ datasets from various domains to train our model. The total number of collected frames is around $9$ million. Details of these datasets and their mixing ratios can be found in the suppl. materials.
Since the quality of data labels are varied, we apply different loss function combinations tailored for different data sources. As shown in Table 1, higher quality and less noisy data are assigned with finer-level loss functions. To address imperfections and potential outliers in real labels, we exclude the top $5\%$ of per-pixel loss values for real data. Following DINOv2 [40], we rebalance the collected datasets using image retrieval from a curated image set to address uneven or biased data distributions. See the suppl. materials for details.
Table 1: Label quality and applied losses for different training sources.
Source Label quality Applied losses
Accuracy Range Density $\mathcal{L}_{G}$ $\mathcal{L}_{S_{1}}$ $\mathcal{L}_{S_{2}}$ $\mathcal{L}_{S_{3}}$ $\mathcal{L}_{N}$ $\mathcal{L}_{M}$
Synthetic Perfect $\infty$ Dense
SfM/MV Recon High $\infty$ Dense&Partial
LiDAR/Laser High $\sim 100$m Sparse
Kinect Medium $\sim 10$m Dense
4 Experiments
Method NYUv2 KITTI ETH3D iBims-1 GSO Sintel DDAD DIODE Average
[…]
Scale-invariant point map
LeReS 17.0 76.0 37.3 11.1 17.1 75.8 18.5 72.1 14.7 75.9 38.6 30.6 31.9 39.6 27.5 46.7 25.3 53.5 3.94
DUSt3R 5.56 97.1 21.9 63.6 10.7 90.5 6.17 95.4 4.55 99.3 34.8 50.4 21.3 70.3 12.3 87.0 14.7 81.7 2.69
UniDepth 5.36 98.3 10.7 98.0 18.5 77.5 5.27 97.4 6.57 99.6 33.0 49.0 11.3 90.3 12.3 91.1 12.9 87.7 2.18
Ours 4.91 98.4 8.62 96.3 4.60 98.9 4.62 97.1 2.60 100 22.3 69.4 12.3 90.4 6.50 94.7 8.32 93.1 1.12
Affine-invariant point map
LeReS 9.55 91.4 29.0 43.8 14.7 79.6 11.0 88.7 8.89 95.3 29.7 55.5 29.3 46.8 14.9 80.3 18.4 72.7 3.94
DUSt3R 4.49 97.4 18.0 66.7 7.27 94.9 5.02 96.0 3.08 99.5 30.3 56.7 19.6 71.4 8.86 88.9 12.1 83.9 2.94
UniDepth 3.96 98.4 7.94 97.3 12.2 89.6 4.62 98.0 2.98 99.8 28.5 58.5 10.2 90.6 8.46 91.1 9.86 90.4 1.88
Ours 3.73 98.3 7.69 95.0 3.60 99.0 3.60 97.3 1.16 100 16.9 77.8 10.4 91.5 4.31 96.6 6.43 94.4 1.25
Local point map
LeReS - - 8.62 91.5 15.7 84.4 - 18.7 85.8 1.65 99.0 16.2 85.2 12.2 89.2 3.70
DUSt3R - - 5.88 94.4 9.43 91.3 - 12.7 88.3 1.71 98.9 10.1 91.3 7.96 92.9 2.30
UniDepth - - 7.45 92.7 10.2 90.2 - 17.0 86.9 1.64 98.8 13.2 87.8 9.89 91.3 3.00
Ours - - 2.86 97.9 7.02 94.1 - 9.73 92.1 1.08 99.5 5.96 95.2 5.33 95.8 1.00
[…]
Method NYUv2 KITTI ETH3D iBims-1 GSO Sintel DDAD DIODE Average
[…]
Scale-invariant depth
LeReS 12.1 82.5 19.3 64.3 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.1 18.1 69.9 18.5 69.5 7.31
ZoeDepth 5.65 96.3 7.23 92.2 10.4 87.3 7.48 93.2 3.23 99.9 27.4 61.7 17.0 73.0 11.2 85.4 11.2 86.1 5.50
DUSt3R 4.43 97.1 7.71 90.9 6.07 95.6 4.96 95.9 3.27 99.5 31.1 57.2 18.5 73.5 8.84 88.8 10.6 87.3 5.00
Metric3D V2 4.74 97.3 3.96 98.6 3.92 98.5 4.20 97.7 2.47 99.9 20.7 70.0 7.36 94.6 3.27 98.5 6.32 94.4 2.07
UniDepth 3.88 98.3 3.73 98.7 5.68 97.0 4.76 97.4 4.17 99.7 28.3 58.9 9.98 90.7 6.79 92.9 8.41 91.7 3.00
DA V1 4.81 97.5 5.52 95.8 8.80 92.3 5.54 95.8 5.48 99.2 28.1 56.7 13.1 81.7 10.1 87.8 10.2 88.4 5.67
DA V2 5.09 97.2 7.28 93.6 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 14.6 78.2 7.89 90.3 9.08 89.7 4.06
Ours 3.48 98.4 4.27 97.8 3.38 98.9 3.45 97.1 1.49 100 19.4 73.4 9.11 90.6 4.85 94.8 6.17 93.9 1.62
Affine-invariant depth
Marigold 4.63 97.3 7.29 93.8 6.08 96.3 4.35 97.2 2.78 99.9 21.2 75.0 14.6 80.5 6.34 94.3 8.41 91.8 2.25
GeoWizard 4.69 97.4 8.14 92.5 6.90 94.0 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.44 90.7 2.69
Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.56
Affine-invariant disparity
MiDaS V3.1 4.62 98.1 6.30 94.8 5.80 96.8 4.75 97.4 1.88 100 21.4 73.0 14.4 82.7 6.06 94.9 8.15 92.2 3.69
DA V1 4.24 98.4 5.40 97.0 4.70 98.2 4.20 97.6 1.54 100 20.4 77.5 12.6 87.0 5.72 95.7 7.36 93.9 2.31
DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 2.50
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.944 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.06
Table 3: Quantitative results for depth map estimation. Gray numbers denote models trained on respective benchmarks and thus excluded from ranking.
Method NYUv2 ETH3D iBims-1 Average
[…]
Perspective 5.38 4.39 13.6 11.9 10.6 9.30 9.86 8.53 5.00
WildCam 3.82 3.20 7.70 5.81 9.48 9.08 7.00 6.03 3.00
LeReS 19.4 19.6 8.26 7.19 18.4 17.5 15.4 14.8 5.53
DUSt3R 2.57 1.86 5.77 3.60 3.83 2.53 4.06 2.66 1.67
UniDepth 7.56 4.31 10.7 9.96 11.9 5.96 10.1 6.74 4.50
Ours 3.41 3.21 2.50 1.54 2.81 1.89 2.91 2.21 1.50
Table 4: Evaluation results for camera FoV in degrees.
Ablation Point Depth Disparity
Scale-inv. Affine-inv. Local Scale-inv. Affine-inv. Affine-inv.
[…]
Depth+camera 11.3 88.7 9.20 90.4 10.1 91.2 8.97 90.1 7.29 92.6 8.22 92.1
Opt. scale-inv. 10.5 89.7 8.50 91.4 9.09 92.0 8.28 90.9 6.72 93.3 7.91 92.6
Opt. affine-inv. 10.1 90.3 8.07 91.9 7.84 93.2 7.92 91.2 6.29 93.7 7.43 93.2
Med. affine-inv. 11.5 88.5 9.36 89.7 10.2 90.9 9.11 89.7 7.38 92.4 8.53 92.0
Full w/o clip 10.1 90.4 8.13 91.4 6.94 94.0 7.94 91.4 6.31 93.5 7.45 93.1
Full w/o $\mathcal{L}_{S}$ 10.1 90.2 8.10 91.8 7.51 93.5 7.94 91.2 6.31 93.6 7.47 93.2
Full 10.0 90.5 8.00 91.8 7.07 94.0 7.84 91.4 6.21 93.8 7.31 93.3
Table 5: Quantitative ablation study results. All experiments are conducted with a ViT-Base backbone for the encoder.
Figure 6: Qualitative ablation study results. All experiments use the ViT-Base backbone for the encoder. (a) Surface visualization of the impact of $\mathcal{L}_{S}$ and $\mathcal{L}_{N}$. Removing either leads to noisy surfaces and poor geometry. (b) Depth visualization for the ablation of valid region mask prediction. Our method correctly predicts the sky regions, for which the predicted point values will be erroneous if the masks are removed. Supervising infinity regions by assigning a large distance label can negatively affect the foreground prediction accuracy.
Implementation details.
We employ a ViT [15] encoder and a light-weight CNN-based upsampler as the decoder. The encoder is pretrained with DINOv2 [40]. The initial learning rates for the encoder and decoder are set to $5\times 10^{-6}$ and $5\times 10^{-5}$, with a decay factor of $5$ applied for every 100K iterations. The training batch size is $256$. Our model is trained on images with varying aspect ratio and resolutions, where aspect ratios are uniformly sampled between $1:2$ and $2:1$, and pixel counts range from 250K to 500K. We apply several image augmentations, including color jittering, Gaussian blurring, JPEG compression-decompression, and random cropping. The principal point of the cropped image is aligned with its center using a perspective crop.
4.1 Quantitative Evaluations
We assess the zero-shot performance of our model and compare it with several state-of-the-art methods for monocular point map estimation [69, 41, 57] and depth map estimation [65, 66, 6, 26, 5, 20, 29]. To ensure a fair comparison, all methods utilize ViT-Large [15] as backbone, except for LeReS (which employs ResNeXt101 [63]) as well as GeoWizard and Marigold (both use StableDiffusion V1 [45]). For a valid evaluation, the scale and shift of the predicted point maps and depth maps are first aligned with the ground truth when necessary. We also evaluate the FoV prediction results against several recent methods [28, 75]. As detailed below, our method achieves superior results across all evaluation metrics for various tasks on different benchmarks.
Point map estimation.
We evaluate the accuracy of point map estimation on eight diverse datasets: NYUv2 [38], KITTI [51], ETH3D [49], iBims-1 [31], Sintel [9], Google Scanned Objects (GSO) [16], DDAD [24], and DIODE [52]. These datasets represent a wide range of domains, including indoor scenes, street views, object scans, and synthetic movies. The raw datasets are processed for reliable evaluation (e.g., sky regions in Sintel dataset and boundary artifacts in the DIODE are removed; see the suppl. material for more preprocessing details). For evaluation metrics, we use the relative point error Relp, i.e., $\|\mathbf{p}-\hat{\mathbf{p}}\|_{2}/\|\hat{\mathbf{p}}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\mathbf{p}-\hat{\mathbf{p}}\|_{2}/\|\hat{\mathbf{p}}\|_{2}<0.25$.
We evaluate our method and compare it with LeReS [69], UniDepth [41], and DUSt3R [57]. LeReS and UniDepth predict depth and camera parameters from single images. DUSt3R is designed for two-view input and directly outputs camera-space point maps. It accepts two identical images for handling monocular scenarios. We evaluate both scale-invariant point maps in camera space and affine-invariant point maps. Table 2 demonstrates that our method outperforms existing methods across most benchmarks, achieving the lowest average Relp for both point maps representations.
To further evaluate local geometry accuracy, we assess affine-invariant point maps within local regions defined by object segmentation masks, either from the dataset labels or obtained using Segment Anything [30]. Since per-region point map evaluation demands high-quality ground truth labels as well as the composition of multiple objects in a scene, NYUv2, KITTI, and GSO are excluded. Compared to the previous methods, our method significantly reduces the region-wise Relp by a large margin: from $7.96$ to $5.33$, achieving approximately a $30\%$ error reduction.
Depth map estimation.
We employ the same benchmarks used for point map evaluation to assess the depth estimation accuracy. We adopt absolute relative error Reld, i.e., $|z-\hat{z}|/z$, and percentage of inliers $\delta_{1}^{d}$ with $\max(z/\hat{z},\hat{z}/z)<1.25$ as the evaluation metrics.
To conduct a comprehensive comparison with different monocular depth estimation methods, we evaluate scale-invariant depth [69, 5, 57, 26, 41], affine-invariant depth [29, 20], and affine-invariant disparity [6, 65, 66]. The $z$ coordinates in our predicted point maps represent affine-invariant depth. By utilizing the camera shifts derived from Eq. 2, we convert these coordinates into scale-invariant depth. For affine-invariant disparity, we take the inverse of our scale-invariant depth predictions.
As shown in Table 3, our method achieves the best performance across most benchmarks and exhibits the lowest average Reld for all depth representations. Since existing methods with scale-invariant depth representations can be converted to affine-invariant depth and disparity, we provide an extensive comparison in the suppl. materials, where our method also significantly outperforms the others.
Camera FoV estimation.
To evaluate the accuracy of the estimated camera FoV, we select benchmarks with vertical $\text{vertical FoV}\geq 45^{\circ}$, including NYUv2 [38], ETH3D [49] and iBims-1 [31], and augment the FoV by randomly center cropping 50$\sim$100from the original images. All input images are undistorted and have a centered principal point.
We compare our model to point map estimation methods [69, 41, 57] as well as two learning-based camera calibration methods: Perspective [28] and WildCam [75]. For fair assessments of zero-shot performance, we utilize their pre-trained checkpoints on in-the-wild datasets. Mean and median FoV errors are reported in Table 5. Our method achieves a mean error of $2.91^{\circ}$ and a median error of $2.21^{\circ}$ on average of the three FoV benchmarks, outperforming both camera calibration methods and point map estimation methods.
4.2 Qualitative comparison
We show visual results of disparity and point maps of different methods in Fig. 7. For depth prediction methods that do not recover camera parameters, we lift their results to 3D points using ground-truth intrinsics when available. Upon visual inspection, our method produces significantly less-distorted point maps compared to others, demonstrating its superior generalizabiltiy and accuracy. For example, as shown in the fourth row of Fig. 7, other methods result in noticeable stretching and deformation of the car shape. Although [66] produces visually sharp disparity maps, distorted 3D geometry can be observed even when provided with ground-truth camera intrinsics. In contrast, our method yields a more regular and well-formed car structure.
Figure 7: Qualitative comparison of point and disparity maps. *: for methods without camera intrinsics prediction, we use ground-truth camera intrinsics or disparity shifts to lift their results into 3D points when available. The first four cases are samples from our test datasets, while the last two are random “in-the-wild” images.
4.3 Ablation Study
We conducted comprehensive ablation studies on our point map representation, alignment strategy and loss function designs. For efficiency, we use the ViT-base model as backbone for these experiments.
Point map representation.
Our method predicts affine-invariant point maps. Here we compare this representation with two variants, one scale-invariant point map in camera space, and another that predicts a depth maps and camera parameters. For the former, we apply the proposed ROE alignment solver to solve the scale parameter only. For the latter, we adopt the SI-log loss [18] for the depth map and the MSE loss for the ray direction map, similar to UniDepth [41]. The first three rows of Table 5 show the results of different approaches. In these experiments, only the global ground-truth supervision is applied during training. Our proposed affine-invariant representation outperforms the other variants across various tasks. This can be attributed to the fact that the other two variants suffer from focal-distance ambiguity issue during training which jeopardizes their performance.
Optimal alignment.
To evaluate the effectiveness of our optimal alignment, we compare it to a previously-used simple alignment method [42], i.e., we normalize the predicted and ground-truth point maps by their median $z$: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})$ equals to $(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})$ is computed as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$. As shown in the third and fourth rows of Table 5, our proposed alignment method consistently outperforms the median alignment strategy across all tasks by large margins, demonstrating the critical role of our proposed optimal alignment.
Multi-scale local geometry loss.
As shown in the six-th row of Table 5, the quantitative results without $\mathcal{L}_{S}$ suffer a significant drop on local geometry metrics. This demonstrates the effectiveness of $\mathcal{L}_{S}$ in enhancing local geometry accuracy and compensating for the insufficient supervision of the global loss due to the ambiguous relative positions between distant scene objects. Fig. 6 (a) further demonstrates the effectiveness of the proposed local geometry loss.
Infinity mask.
Our method predicts a valid region mask to handle regions at infinity. As shown in Fig. 6 (b), our method correctly predicts the sky regions, for which the predicted point values will be erroneous if the masks are removed. We further compare our mask prediction design with another strategy for addressing infinity. Specifically, we assign a large value, i.e., $1,000$ times the average distance of valid regions, to the infinity regions. Fig. 6 (b) shows that this strategy compromises the accuracy of foreground predictions as the network struggles to generate large values for infinity. Incorporating a separate mask prediction offers a convenient solution for masking out infinity and retaining accurate foreground geometry learning.
5 Conclusion
We have presented a method for accurate monocular geometry estimation of open-domain images. Our key insight is to design effective supervision which is largely neglected by previous methods. We propose a new affine-invariant point map representation which precludes ambiguous supervision during training. A robust, optimal, and efficient point map alignment solver is introduced for accurate global shape learning, together with a multi-scale local geometry loss designed for precise local geometry supervision. Trained on a large dataset collection, our model demonstrates strong generalizability and high accuracy for open-domain images. It significantly outperforms previous methods across various tasks and benchmarks. We believe our method marks a significant advancement in monocular 3D geometry estimation and can serve as a robust foundation model for various vision applications.
Supplementary Material
Appendix A Algorithm details
A.1 Recovering shift and camera focal
We assume a simple pinhole camera model with isotropic focal length and centered principal point. The 2D image plane is parameterized with the center as $(0,0)$. The image plane coordinate of pixel $i$ is denoted as $(u_{i},v_{i})$, corresponding to its predicted 3D point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$. The focal length and shift is obtained by minimizing the projection error,
$\min_{f,t_{z}}\sum_{i\in\mathcal{M}}\left({fx_{i}\over z_{i}+t_{z}^{\prime}}-u_{i}\right)^{2}+\left({fy_{i}\over z_{i}+t_{z}^{\prime}}-v_{i}\right)^{2}$ (9)
which can be further reduced to have a single variable $t_{z}^{\prime}$ by substituting $f$ with its close-form solution with respect to $t_{z}^{\prime}$
$f={\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}\right)u_{i}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}\right)v_{i}\over\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}\right)^{2}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}\right)^{2}}$ (10)
We use a numerical solver for this least squares problem with Levenberg-Marquardt algorithm [37] implemented by SciPy [53] package. For efficiency, the point map is resized to low resolution ($64\times 64$) for running this algorithm. In our practice, it typically converges within 10 iterations in around 3ms.
A.2 Optimal alignment
We will first introduce an algorithm to a simpler subproblem then derive the solution to either with or without the constraint of $t_{x}=t_{y}=0$ (1D-shift case or 3D-shift case, respectively).
Subproblem (w/o clipping)
Consider the optimization objective with respect to scale $s$ only, denoted as $l_{0}(s)$. We omit the mask $\mathcal{M}$ for simplicity and denote $N$ as the number of valid points.
$\min_{s}l_{0}(s)=\min_{s}\sum_{i=1}^{N}w_{i}|s\hat{x}_{i}-x_{i}|$ (11)
where $w_{i}>0$ and $\hat{x}_{i}>0$ without loss of generality. The objective, as a summation of convex functions, is also convex obviously. The minimum occurs where its left-hand $l_{0}^{\prime-}(s)$ derivative and right-hand derivative $l_{0}^{\prime+}(s)$ have opposite signs or one of them is zero,
$\begin{split}l_{0}^{\prime-}(s)&=\sum_{{x_{i}\over\hat{x}_{i}}<s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\sum_{i=1}^{N}w_{i}\hat{x}_{i}\\ l_{0}^{\prime+}(s)&=\sum_{{x_{i}\over\hat{x}_{i}}\leq s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\sum_{i=1}^{N}w_{i}\hat{x}_{i}\end{split}$ (12)
$l_{0}^{\prime-}(s)$ and $l_{0}^{\prime+}(s)$ differ at $\left\{x_{i}\over\hat{x}_{i}\right\}$. One might first sort $\left\{x_{i}\over\hat{x}_{i}\right\}$ and compute the prefix summations of $\{w_{i}\hat{x}_{i}\}$. Then the derivatives can be evaluated in $O(1)$ time for each point in $\left\{x_{i}\over\hat{x}_{i}\right\}$. Finally, $\hat{x}_{i}\over x_{i}$ such hat ł$\l_{0}^{\prime-}\left(x_{i}\over\hat{x}_{i}\right)<0\leq\l_{0}^{\prime+}\left(x_{i}\over\hat{x}_{i}\right)$ is a deterministic minimum point.
Subproblem (w/ clipping)
We clip each residual term to suppress the contribution of outliers. Specifically, the clipped objective is
$\min_{s}l_{1}(s)=\min_{s}\sum_{i=1}^{N}\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ (13)
For each item $l_{1,i}(s)=\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ in the equation, the one-sided derivatives are
$\begin{split}l_{1,i}^{\prime-}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.\\ l_{1,i}^{\prime+}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.\end{split}$ (14)
Therefore, the one-sided derivatives of $l_{0}(s)$ are
$\begin{split}l_{1}^{\prime-}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime-}(s)\\ &=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}\end{split}$ (15)
$\begin{split}l_{1}^{\prime+}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime+}(s)\\ &=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}\end{split}$ (16)
One can evaluate $l_{1}^{\prime}(s)$ efficiently through first binary searching on the sorted arrays $\left\{x_{i}\over\hat{x}_{i}\right\}$, $\left\{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}\right\}$ and $\left\{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}\right\}$ and then indexing the precomputed prefix summations of $\{w_{i}\hat{x}_{i}\}$ in the associated orders.
Lemma 1.
There exists at least one pair of $(k^{*},s^{*})$ such that $s^{*}x_{k^{*}}-\hat{x}_{k^{*}}=0$ and $s^{*}$ minimizes Eq. 13
Proof.
The minimum of $l_{1}(s)$ must exist, because $l_{1}(s)$ is continuous, piece-wisely linear and bounded in $[0,N\tau]$.
We first prove that there must exist $s^{*}$ such that $l_{1}(s^{*})=\min l_{1}(s)$ and $l_{1}^{\prime+}(s)>l_{1}^{\prime-}(s)$. Otherwise, for all $s^{*}$ such that $l_{0}(s^{*})=\min l_{1}(s)$, there will be $l_{1}^{\prime+}(s)=l_{1}^{-}(s)=0$, hence the value of $l_{1}(s)$ in the linear interval where the minimum locates is constant. As a consequence, all neighboring intervals will be constant until the boundary where $\min l_{1}(s)=l_{1}(-\infty)=N\tau$, which contradicts the obvious fact that $\min l_{1}(s)\leq l_{1}(x_{1}/\hat{x}_{1})<N\tau$.
Given $l_{0}^{\prime+}(s^{*})>l_{0}^{\prime-}(s^{*})$, there exists an index $k^{*}$ such that $s^{*}=\hat{x}_{k^{*}}/x_{k^{*}}$, because
$\begin{split}0<&l_{1}^{\prime+}(s^{*})-l_{1}^{\prime-}(s^{*})\\ =&2\sum_{{\hat{x}_{i}\over x_{i}}=s^{*}}w_{i}x_{i}-\!\!\!\!\!\!\sum_{{w_{i}\hat{x}_{i}-\tau\over w_{i}x_{i}}=s^{*}}\!\!\!\!w_{i}x_{i}-\!\!\!\!\!\!\sum_{{w_{i}\hat{x}_{i}+\tau\over w_{i}x_{i}}=s^{*}}\!\!\!\!w_{i}x_{i}\\ <&2\sum_{{\hat{x}_{i}\over x_{i}}=s^{*}}w_{i}x_{i}\end{split}$ (17)
Lemma 1 has proved that the minimum point of the objective function in Eq. 13 still occurs among $\left\{x_{i}\over\hat{x}_{i}\right\}$, although the function is non-convex and may contain local minima. Therefore, it takes two steps to solve the subproblem. First, find all extrema among $\left\{x_{i}\over\hat{x}_{i}\right\}$ such that $l_{0}^{\prime-}\left(x_{i}\over\hat{x}_{i}\right)<0\leq l_{0}^{\prime+}\left(x_{i}\over\hat{x}_{i}\right)$ by evaluating the derivative values, which takes $O(N\log N)$ time. Second, evaluate the objective values of those extrema and then find the minimum, which takes $O(Nn_{e})$ time. $n_{e}$ is the number of extrema, approximating the number of outliers , which is expected to be a small constant in practice.
Alignment with 1D shift
Recall the objective in Eq. 4. Let $w_{i}$ be $1/z_{i}$. We apply clipping to each absolute term and rewrite it as follows
$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-x_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|)\end{split}$ (18)
Lemma 2.
There exists at least one triplet of $(k^{*},s^{*},t_{z}^{*})$ such that $s^{*}\hat{z}_{k^{*}}+t_{z}^{*}-z_{k^{*}}=0$ and $(s^{*},t_{z}^{*})$ minimizes the objective of Equation 18.
Proof.
Denote the objective as $l_{2}(s,t_{z})$,
$\begin{split}l_{2}(s,t_{z})=&\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{x}_{i}-x_{i}|}\right)\\ &+\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{y}_{i}-y_{i}|}\right)\\ &+\sum_{i=1}^{n}\min\left(\tau,w_{i}{|t_{z}-(z_{i}-s\hat{z}_{i})|}\right)\end{split}$ (19)
Given arbitrary $s$, using Lemma 1, there exists at least one pair $(t_{z},k)$ such that $t_{z}-(z_{k}-s\hat{z}_{k})=0$ and $t_{z}$ minimizes $\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{z}_{i}+t_{z}-z_{i}|}\right)$, hence minimizes $l(s,t_{z})$ as the rest parts are constant with regard to $t_{z}$. Therefore, a solution $s^{*}$ is always associated with corresponding $(t^{*}_{z},k^{*})$ such that that $s^{*}\hat{z}_{k^{*}}+t^{*}_{z}-z_{k^{*}}=0$. ∎
Lemma 2 allows us to reduce Eq. 18 to the subproblem with respect to some index $k$. For each possible index $k$, the objective is formed as
$\begin{split}\min_{s}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-\!sx_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s(\hat{z}_{i}-\hat{z}_{k})-(z_{i}-z_{k})|)\end{split}$ (20)
which is solvable in $O(N\log N)$ complexity. We enumerate all possible indices for $k$ and find the minimum. Therefore, the total time complexity is $O(N^{2}\log N)$.
In our implementation, the point map is resized to low resolution ($64\times 64$) for alignment, with $N=4096$ at most. The algorithm is further parallelized with tensor operations on GPUs.
Alignment with 3D shift
Similarly to the proof of Lemma 2, there exists at least one group $(k_{1}^{*},k_{2}^{*},k_{3}^{*},s^{*},\mathbf{t}^{*})$ such that $s^{*}\hat{x}_{k_{1}^{*}}+t_{x}^{*}-x_{k_{1}^{*}}=0$, $s^{*}\hat{y}_{k_{2}^{*}}+t_{y}^{*}-y_{k_{2}^{*}}=0$, $s^{*}\hat{z}_{k_{3}^{*}}+t_{z}^{*}-z_{k_{3}^{*}}=0$, and $(s^{*},\mathbf{t}^{*})$ minimizes the objective. However, the $O(N^{4}\log N)$ time complexity of a brute-force search is prohibitive. Motivated by the strong locality of surface points within a 3D sphere, we introduce a reasonable assumption, $k_{1}=k_{2}=k_{3}$, to obtain an approximately optimal solution with $O(N^{2}\log N)$ complexity. This assumption posits that the predicted and ground truth patches can be well aligned under the condition that one corresponding pair of points coincides. The effectiveness of the approximated solution has been empirically validated.
Appendix B Data details
B.1 Training data
The datasets used in our training are listed in Table 6. The number of frames may slightly differ from that of the original data because some invalid frames are dropped.
Name Domain #Frames Type Weight
Training
A2D2[21] Outdoor/Driving $196$K C 0.8%
Argoverse2[60] Outdoor/Driving $1.1$M C 7.4%
ARKitScenes[3] Indoor $449$K B 8.6%
DIML-indoor[14] Indoor $894$K D 4.8%
BlendedMVS[67] In-the-wild $115$K B 12.0%
MegaDepth[34] Outdoor/In-the-wild $92$K B 5.6%
Taskonomy[72] Indoor $3.6$M B 14.1%
Waymo[50] Outdoor/Driving $788$K C 6.4%
GTA-SfM[55] Outdoor/In-the-wild $19$K A 2.8%
Hypersim[44] Indoor $75$K A 5.0%
IRS[56] Indoor $101$K A 5.6%
KenBurns[39] In-the-wild $76$K A 1.6%
MatrixCity[33] Outdoor/Driving $390$K A 1.3%
MidAir[19] Outdoor/In-the-wild $423$K A 4.0%
MVS-Synth[27] Outdoor/Driving $12$K A 1.2%
Spring[36] In-the-wild $5$K A 0.7%
Structured3D[74] Indoor $77$K A 4.8%
Synthia[46] Outdoor/Driving $96$K A 1.2%
TartanAir[58] In-the-wild $306$K A 5.0%
UrbanSyn[25] Outdoor/Driving $7$K A 2.1%
ObjaverseV1[12] Object $167$K A 4.8%
Table 6: A table of datasets used for training. #Frames stands for the number of used frames. The types are in short for A. synthetic, B. structure from motion / multiview reconstruction, C. LiDAR / Laser; D. Microsoft Kinect
B.2 Evaluation data
The raw evaluation datasets are processed accordingly for reliable evaluation and fair comparison. We report the details as follows.
  • NYUv2 [38] We use the official test split of 654 samples Due to the inaccuracy of ground truth values captured by Kinect V1 near boundaries, we filter and remove boundary regions by a simple edge detection method. Specifically, we set a threshold for the difference between the minimum and maximum depth values within a local window. Depth values beyond 5 meters are excluded because they are unreliable due to the limited sensor range. Additionally, we manually mask out areas with reflective and transparent surfaces, such as mirrors and glass, which cannot be accurately captured by the sensor.
  • KITTI [51] We utilizes the test split of 652 images of Eigen et. al. [18] following previous works. The original resolution of $1242\times 375$ does not match our training aspect ratio (ranging from $1:2$ to $2:1$), so we apply a center crop to obtain a $750\times 375$ resolution from the raw images.
  • ETH3D [49] All 454 images are included. The images are undistorted with the official calibration data and downsized from the original $6202\times 4135$ resolution to $2048\times 1365$.
  • iBims-1 [31] All 100 images are included at an original resolution of $640\times 480$.
  • GSO [16] The dataset contains 1,030 objects. For each object, we render a single view at $512\times 512$ resolution. The view is randomly sampled with a FoV ranging from $30^{\circ}$ to $60^{\circ}$. The object is centered in the image, and its bounding box occupies approximately 70% of the image’s size.
  • Sintel [9] We use all 1,064 frames and perform a center crop to $872\times 436$ from the original $1024\times 436$ resolution to fit our aspect ratio range. The sky regions are manually masked out.
  • DDAD [24] We randomly select 1,000 samples from the validation set. The dataset was collected using multiple cameras and LiDAR sensors mounted on a moving vehicle. Some cameras inadvertently capture parts of the vehicle, causing discrepancies with the sensor’s depth data. To address this, we crop the regions that are not obstructed by the vehicle itself.
  • DIODE [52] We utilize the official validation split, which includes 325 indoor images and 446 outdoor images at an original resolution of $1024\times 768$. Due to artifacts in ground truth depth values near the boundaries in this dataset, we identify and remove these boundary regions using a similar approach as described above.
Appendix C More results
C.1 Full table of depth estimation results
In Table 7, we present a full table of results for depth estimation where methods that predict metric or scale-invariant depth are evaluated for affine-invariant depth and disparity for fair comparison with other baselines.
Method NYUv2 KITTI ETH3D iBims-1 GSO Sintel DDAD DIODE Average
[…]
Scale-invariant depth
LeReS 12.1 82.5 19.3 64.3 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.1 18.1 69.9 18.5 69.5 7.31
ZoeDepth 5.65 96.3 7.23 92.2 10.4 87.3 7.48 93.2 3.23 99.9 27.4 61.7 17.0 73.0 11.2 85.4 11.2 86.1 5.50
DUSt3R 4.43 97.1 7.71 90.9 6.07 95.6 4.96 95.9 3.27 99.5 31.1 57.2 18.5 73.5 8.84 88.8 10.6 87.3 5.00
Metric3D V2 4.74 97.3 3.96 98.6 3.92 98.5 4.20 97.7 2.47 99.9 20.7 70.0 7.36 94.6 3.27 98.5 6.32 94.4 2.07
UniDepth 3.88 98.3 3.73 98.7 5.68 97.0 4.76 97.4 4.17 99.7 28.3 58.9 9.98 90.7 6.79 92.9 8.41 91.7 3.00
DA V1* 4.81 97.5 5.52 95.8 8.80 92.3 5.54 95.8 5.48 99.2 28.1 56.7 13.1 81.7 10.1 87.8 10.2 88.4 5.67
 - metric indoor 4.81 97.5 15.2 74.2 9.42 89.0 5.54 95.8 5.48 99.2 28.3 56.7 24.2 57.5 10.1 87.8 12.9 82.2 -
 - metric outdoor 16.0 72.2 5.52 95.8 8.80 92.3 13.8 78.8 8.60 93.6 28.1 54.8 13.1 81.7 12.9 81.7 13.3 81.4 -
DA V2* 5.09 97.2 7.28 93.6 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 14.6 78.2 7.89 90.3 9.08 89.7 4.06
 - metric indoor 5.09 97.2 7.49 91.3 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 16.5 73.6 7.89 90.3 9.35 88.8 -
 - metric outdoor 15.3 72.1 7.28 93.6 9.33 89.6 10.6 84.9 9.61 92.5 28.6 57.4 14.6 78.2 12.1 83.6 13.4 81.5 -
Ours 3.48 98.4 4.27 97.8 3.38 98.9 3.45 97.1 1.49 100 19.4 73.4 9.11 90.6 4.85 94.8 6.17 93.9 1.62
Affine-invariant depth
LeReS 6.25 95.4 8.15 90.6 8.98 90.7 6.71 94.6 4.03 99.4 24.0 64.8 16.1 75.9 9.93 88.4 10.5 87.5 8.81
ZoeDepth 4.79 97.2 5.51 95.3 7.29 94.2 5.88 95.7 2.54 99.9 21.8 69.2 14.1 80.3 7.72 91.0 8.70 90.3 7.25
DUSt3R 3.76 97.8 7.21 91.8 4.96 96.4 3.92 96.6 2.56 99.6 25.4 64.1 16.8 76.4 6.55 92.9 8.89 89.5 6.69
Metric3D V2 4.00 97.6 3.48 98.5 3.54 98.4 3.25 98.3 2.12 99.4 24.4 71.6 7.10 94.8 2.73 98.8 6.33 94.7 3.50
UniDepth 3.42 98.6 3.54 98.8 4.93 97.5 3.74 98.2 2.47 99.9 24.8 64.1 9.36 91.0 4.88 96.2 7.14 93.0 3.62
Marigold 4.63 97.3 7.29 93.8 6.08 96.3 4.35 97.2 2.78 99.9 21.2 75.0 14.6 80.5 6.34 94.3 8.41 91.8 5.81
GeoWizard* 4.69 97.4 8.14 92.5 6.90 94.0 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.44 90.7 6.56
- indoor 4.69 97.4 8.14 92.5 6.90 93.9 4.50 97.1 2.00 99.9 17.8 76.0 16.5 75.7 7.03 92.7 8.44 90.7 -
- outdoor 4.83 97.3 8.40 92.2 7.18 93.7 4.59 97.0 2.03 99.9 17.8 76.2 17.1 74.4 7.22 92.5 8.64 90.4 -
- object 4.70 97.4 8.21 92.4 6.92 94.0 4.52 97.0 2.05 99.9 17.9 76.0 16.8 75.1 7.13 92.6 8.52 90.5 -
DA V1* 3.86 98.3 4.94 96.6 6.26 96.5 4.24 97.3 1.98 100 20.1 71.8 11.2 86.2 6.50 93.4 7.55 92.8 4.50
 - metric indoor 3.86 98.3 9.78 87.2 6.35 95.0 4.24 97.3 1.98 100 20.1 71.8 16.9 74.2 6.50 93.0 8.72 89.6 -
 - metric outdoor 7.72 93.8 4.94 96.6 6.26 96.5 7.01 94.2 2.78 99.8 20.7 70.0 11.2 86.2 6.93 93.4 8.44 91.3 -
DA V2* 4.22 97.8 6.81 94.2 4.66 97.2 3.45 98.3 1.45 100 17.2 76.5 14.2 82.0 5.32 94.9 7.16 92.6 3.19
 - metric indoor 4.22 97.8 6.95 92.7 4.66 97.2 3.45 98.3 1.45 100 17.2 76.5 14.2 80.0 5.32 94.9 7.18 92.2 -
 - metric outdoor 8.69 90.9 6.81 94.2 7.26 93.5 6.81 93.5 2.28 100 22.5 67.0 13.2 82.0 8.00 91.2 9.45 89.0 -
Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.56
Affine-invariant disparity
LeReS 7.33 95.4 12.4 86.7 10.2 90.1 8.45 92.9 4.33 99.7 28.9 59.6 23.3 73.1 10.8 88.3 13.2 85.7 8.25
ZoeDepth 5.24 97.7 5.86 95.6 8.08 94.0 6.22 96.1 2.59 99.9 26.9 66.3 14.0 81.8 8.17 92.0 9.63 90.4 6.75
DUSt3R 4.27 98.1 7.67 92.2 5.60 96.2 4.47 96.7 2.64 99.8 39.9 56.8 17.4 76.3 7.09 92.8 11.1 88.6 6.75
Metric3D V2 13.4 81.4 3.69 98.3 4.38 97.6 8.56 92.3 1.81 100 21.8 72.5 7.32 94.2 7.62 90.4 8.57 90.8 5.29
UniDepth 3.80 98.7 3.64 98.7 5.34 97.2 4.04 98.1 2.55 99.9 28.5 60.9 9.83 89.3 5.94 95.5 7.96 92.3 3.62
MiDaS V3.1 4.62 98.1 6.30 94.8 5.80 96.8 4.75 97.4 1.88 100 21.4 73.0 14.4 82.7 6.06 94.9 8.15 92.2 5.00
DA V1 4.24 98.4 5.40 97.0 4.70 98.2 4.20 97.6 1.54 100 20.4 77.5 12.6 87.0 5.72 95.7 7.36 93.9 3.00
DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 3.06
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.944 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.38
Table 7: Full table of comparison for depth map estimation. * Methods have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results in smaller text size for each version. Gray numbers denote models trained on respective benchmarks.
C.2 Qualitative comparison on uncurated in-the-wild images
We further utilize the first 100 in-the-wild images in DIV2K[2] dataset to extensively compare our method with LeReS [69], UniDepth [41], DUSt3R [57], Metric3D V2 [26] and Depth Anything V2 [66]. The results can be found in the supplementary videos. Since Metric3D V2 and Depth Anything V2 predict depth map and require ground truth camera focal to obtain 3D points cloud results, we visualize them using our estimated focal lengths.
C.3 More visual results
In the following pages, we demonstrate more examples of our results for in-the-wild images. The rows from left to right are the input images, the disparity maps, surface geometry visualization from the source view and three novel view image of the point cloud, respectively.

### v2

@@
 Given a single image, our model directly predicts a 3D point map of the captured scene with an affine-invariant representation, which is agnostic to true global scale and shift.
-This new representation precludes ambiguous supervision in training and facilitate effective geometry learning.
-Furthermore, we propose a set of novel global and local geometry supervisions that empower the model to learn high-quality geometry.
+This new representation precludes ambiguous supervision in training and facilitates effective geometry learning.
+Furthermore, we propose a set of novel global and local geometry supervision techniques that empower the model to learn high-quality geometry.
 These include a robust, optimal, and efficient point cloud alignment solver for accurate global shape learning, and a multi-scale local geometry loss promoting precise local geometry supervision.
@@
 In our comprehensive evaluation on diverse unseen datasets, our model significantly outperforms state-of-the-art methods across all tasks, including monocular estimation of 3D point map, depth map, and camera field of view.
-Code and models will be released on our project page.
+Code and models can be found on our project page.
 {strip}[Uncaptioned image]
@@
 The model yields high-quality shapes and generalizes well to open-domain images.
-(Best view with zoom)
+(Best viewed with zoom)
 1 Introduction
@@
 For monocular geometry estimation (MGE), a common approach involves first estimating a depth map, subject to an unknown scale (and shift), and then combining it with camera intrinsics to recover the 3D shape via unprojection [71, 42].
-For the former, significant advancements in monocular depth estimation (MDE) have been achieved in recent years by training on large-scale data [71, 65, 29, 20].
+For the former, significant advancements in monocular depth estimation (MDE) have been achieved in recent years by training on large-scale data [73, 26, 67, 68, 43] or leveraging powerful pretrained models [29, 20, 23].
 However, inferring camera intrinsics (e.g., focal length) from single images remains challenging due to the high degree of ambiguity when strong geometric cues are absent.
@@
 More importantly, our findings indicate that the design of training supervisions is crucial to achieve robust and accurate point map prediction from single images.
-Similar to previous MDE approaches, we need to compute a global scaling factor and a translation to align the predicted results with the ground truth during training.
-However, previous methods for this global alignment calculation are either sensitive to outliers or solved with coarse approximations, leading to unsatisfactory supervision.
-We propose a Robust, Optimal, and Efficient (ROE) global alignment solver to resolve scale and shift for the affine-invariant pointmap loss, which substantially improved the training effectiveness and final accuracy.
-On the other hand, the effective learning of local, region-wise geometry has been largely neglected previously.
-In monocular geometry estimation, the relative distance between different objects can be ambiguous, which hinders precise local geometry learning when a global alignment computed with all objects is applied.
-In light of this, we propose a multi-scale local geometry loss, which imposes penalties on the local discrepancies of 3D point clouds, each under independent and optimal affine alignment.
+Similar to previous MDE approaches, a global scaling factor and translation are required to align predictions with ground truth during training.
+However, existing methods for this global alignment calculation are either sensitive to outliers or solved with coarse approximations, leading to suboptimal supervision.
+We propose a Robust, Optimal, and Efficient (ROE) global alignment solver to resolve scale and shift for the affine-invariant point map loss, which substantially improves the training effectiveness and final accuracy.
+On the other hand, the effective learning of local, region-specific geometry has been largely neglected previously.
+In monocular geometry estimation, the relative distances between different objects can be ambiguous, which hinders precise local geometry learning when only a global alignment is applied.
+In light of this, we propose a multi-scale local geometry loss, which penalizes local discrepancies in 3D point clouds under independent affine alignments.
 This design significantly boosted the accuracy of our local geometry prediction.
-We train our model on a large-scale training corpus sourced from various existing datasets.
-Our model demonstrates strong generalizability and accuracy in monocular geometry estimation for open-domain images (some examples are shown in Fig.
+We train our model on large-scale data sourced from various existing datasets.
+Our model demonstrates strong generalizability and accuracy in monocular geometry estimation for open-domain images (examples in Fig.
 1).
-The zero-shot evaluation on eight unseen datasets shows that it outperforms previous MGE methods that output point clouds by a large margin (over 35% error reduction compared to the previous best).
-Moreover, our method also significantly outperforms previous approaches that focus on the sub-tasks of MDE and camera FoV (20$\sim$30error reduction for the former and over 20% for the latter).
-Figure 2 shows the performance ranking of different methods in our evaluation, where our method ranks top across all tasks/metrics and among all compared methods.
+Zero-shot evaluation on eight unseen datasets shows a significant improvement over previous MGE methods that output point clouds (over 35% error reduction compared to the previous best).
+Moreover, our method significantly outperforms prior approaches that focus on the sub-tasks of MDE and camera FOV (20$\sim$30error reduction for the former and over 20% for the latter).
+Figure 2 shows that our method ranks top across all tasks and metrics in the evaluation.
 Our contributions are summarized as follows: 1) We propose a new direct MGE method for open-domain images using affine-invariant point maps.
@@
 3) We demonstrate the substantial enhancement in performance of our model over existing methods on MGE, MDE, and camera FOV estimation across diverse datasets.
-We hope that our method can serve as a strong and versatile foundational model for solving monocular geometry problems, facilitating applications such as 3D-aware image editing, image relighting, depth-to-image synthesis, novel view synthesis, and 3D scene understanding, as well as providing initial geometrical priors for video-based or multiview-based 3D reconstruction.
+We hope that our method serves as a versatile foundational model for monocular geometry tasks, facilitating applications such as 3D-aware image editing, depth-to-image synthesis, novel view synthesis, and 3D scene understanding, as well as providing initial geometrical priors for video or multiview-based 3D reconstruction.
 2 Related Work
@@
 Monocular point map estimation.
-Compared to depth map, 3D point map offers a more straightforward representation of 3D geometry.
-Monocular point map estimation aims to recover free 3D points for each pixel on a single image.
-Several approaches tackle this challenge by predicting camera parameters alongside depth estimation.
-For example, LeReS [69, 70] introduced a two-stage pipeline with an affine-invariant depth prediction module followed by a point cloud module for recovering shift and camera focal length.
+Compared to depth maps, 3D point map offers a more straightforward representation of 3D geometry.
+Monocular point map estimation aims to recover free 3D points for each pixel in a single image.
+Several approaches address this by predicting camera parameters alongside depth.
+For example, LeReS [71, 72] introduced a two-stage pipeline with affine-invariant depth prediction followed by a point cloud module to recover shift and camera focal length.
 UniDepth [42] proposed a self-promptable camera module that predicts dense camera representation to condition the subsequent depth estimation module.
-DUSt3R [57] employs a more straightforward end-to-end model to directly map two-view images to camera space point maps, and it can adapt to monocular scenarios by using two identical images as input.
-It uses the camera-space scale-invariant pointmap which incorporates camera shift information.
-This representation can suffer from the focal-distance ambiguity and jeopardize pointmap learning.
+DUSt3R [59] employs an end-to-end model to directly map two-view images to camera space point maps, adaptable to monocular scenarios by using two identical input images.
+However, its scale-invariant point map can suffer from the focal-distance ambiguity.
 In contrast, our approach is designed to handle monocular input and we employ affine-invariant point map with meticulously designed training supervisions for more effective geometry learning.
@@
 While traditional multiview camera calibration using special patterns is well studied [75], estimating camera intrinsics from single images is ill-posed and remains challenging.
-For the latter, earlier works utilized known 3D shapes [59] or vanishing points [13] to solve the problem.
-More recently, learning-based approaches have been proposed [61, 7, 32, 28, 75] to handle in-the-wild images, but the results are far from satisfactory.
+For the latter, earlier works utilized known 3D shapes [61] or vanishing points [13].
+More recently, learning-based approaches have been explored [63, 7, 33, 28, 77] to handle in-the-wild images, but the results are far from satisfactory.
 In this work, we demonstrate that our camera parameter estimation derived from point maps can achieve state-of-the-art accuracy and offer remarkable generalizability.
@@
 The affine-invariant point map representation can be used to recover both camera shift and focal length.
-Given the predicted 3D points $(x_{i},y_{i},z_{i})$ and their corresponding 2D pixels $(u_{i},v_{i})$, we solve for camera focal length $f$ and $Z$-axis shift by minimizing the projection error:
+Given the predicted 3D points $(x_{i},y_{i},z_{i})$ and their corresponding 2D pixels $(u_{i},v_{i})$, we solve for camera focal length prediction $f$ and $Z$-axis shift $t_{z}^{\prime}$ by minimizing the projection error:
 $\min_{f,t_{z}^{\prime}}\sum_{i=1}^{N}\left({fx_{i}\over z_{i}+t_{z}^{\prime}}-u_{i}\right)^{2}+\left({fy_{i}\over z_{i}+t_{z}^{\prime}}-v_{i}\right)^{2},$ (2)
@@
 Eq.
-2 can be efficiently solved with a fast iterative optimization algorithm, which converges within 10 iterations, taking approximately $3$ms.
+2 can be efficiently solved with a fast iterative optimization algorithm, which typically converges within 10 iterations, taking approximately $3$ms.
 More details about the solution can be found in the suppl.
@@
 Let $\hat{\mathbf{p}}_{i}$ denote the predicted 3D point for the $i$-th pixel, and $\mathbf{p}_{i}$ the corresponding ground truth. The global point map loss is defined as:
-$\mathcal{L}_{G}=\sum_{i\in{\mathcal{M}}}{\frac{1}{{z}_{i}}\|s\hat{\mathbf{p}}_{i}+\mathbf{t}^{*}-{\mathbf{p}}_{i}\|_{1}},$ (3)
+$\mathcal{L}_{G}=\sum_{i\in{\mathcal{M}}}{\frac{1}{{z}_{i}}\|s\hat{\mathbf{p}}_{i}+\mathbf{t}-{\mathbf{p}}_{i}\|_{1}},$ (3)
 where $\|\!\cdot\!\|_{1}$ denotes the $L_{1}$ norm, $s$ and $\mathbf{t}$ are the alignment parameters that transform the predicted affine-invariant point map to the ground-truth camera space, and ${\mathcal{M}}$ is the mask for regions with labels.
@@
 To apply the global loss $\mathcal{L}_{G}$ for training, we need to first determine $s$ and $\mathbf{t}$.
-Previous methods for affine-invariant depth estimation often use coarse approximations for this alignment.
-For example, a predominate strategy is to use the median depth values of the two sets as the anchor to compute shift followed by scaling factor calculation [42, 65, 66].
-The range normalization strategy used by [29, 20] can be viewed as computing shift using minimal or a predefined fraction of depth values.
-All these simple strategies are non-optimal for the problem setup and could lead to unsatisfactory supervision.
-In this work, we propose an optimal solver to the alignment parameters. Specifically, we determine $s^{*}$ and $\mathbf{t}^{*}$ via
+Previous affine-invariant depth estimation methods often use rough alignment approximations, like using the median depth of two sets as an anchor to compute shift, followed by scaling [43, 67, 68], or normalizing depth range [29, 20].
+These simple strategies yield suboptimal alignment and could lead to unsatisfactory supervision.
+In this work, we propose a solver to the optimal alignment parameters. Specifically, we determine $s^{*}$ and $\mathbf{t}^{*}$ via
 $\begin{split}(s^{*},\mathbf{t}^{*})=\mathop{\arg\!\min}_{s,\mathbf{t}}\sum_{i\in{\mathcal{M}}}\frac{1}{{z}_{i}}\left\|s\hat{\mathbf{p}}_{i}+\mathbf{t}-{\mathbf{p}}_{i}\right\|_{1},\end{split}$ (4)
 where $t_{x}$ and $t_{y}$ are 00.
-One possible approach to solve this equation is to formulate it as an absolute residuals optimization problem and solve it using linear programming techniques such as simplex algorithms and the interior-point method [8].
-However, these methods are known to have high computational complexity, often exceeding $O(N^{3})$.
-Given thousands of points, a typical implementation may require computation time of a few seconds or even higher, which is inefficient for network training.
-Instead, we develop an efficient parallelized searching algorithm to solve it optimally.
-Our theoretical analysis (Sec.
-A.2) demonstrates that the minimum must occur when $s\hat{z}^{(k)}+t_{z}-z^{(k)}=0$ for some $k=1,\ldots,N,\ k\in\mathcal{M}$.
-This insight allows us to substitute $t_{z}$ with $s$ and decompose the problem into a series of one-dimensional subproblems with scale variable only, which can be solved in parallel.
-The overall computational complexity is $O(N^{2}\log N)$ and efficient training is achieved on GPU.
-A high-level description of this solution is presented in Algorithm 1 and more details can be found in the suppl.
+One approach to solve this equation is to frame it as an absolute residuals optimization problem and use linear programming methods such as simplex algorithms and the interior-point method [8].
+However, these methods are known to have high computational complexity, often exceeding $O(N^{3})$, making them inefficient for network training with thousands of points, where computation can take seconds.
+Instead, we develop an efficient parallelized searching algorithm.
+With the fact that the optimum must occur when $s\hat{z}_{k}+t_{z}=z_{k}$ for some index $k$, we replace $t_{z}$ with $s$ and break the problem into a series of parallel one-dimensional subproblems.
+This reduces the complexity to $O(N^{2}\log N)$ and enables efficient GPU-based training.
+A high-level description of this solution is presented in Algorithm 1 with more details in the suppl.
 materials.
-In practice, we noticed that Eq.
-4 is still sensitive to outliers even if the $L_{1}$ error norm has been used.
-If the model wrongly predicts a close foreground edge pixel as a far background point, the objective will be dominated by the incorrect edge prediction.
-To further improve the robustness, we apply a clipping operator $\operatorname{clip}\left(\cdot,\tau\right)=\max\big{(}\min(\cdot,\tau),-\tau\big{)}$ on the per-point residuals, which brings accuracy gains as we will show in the experiments.
-Adding the clipping operator makes the problem non-convex, but our optimal solver is still applicable with minor modifications.
+In practice, we found that Eq.
+4 is still sensitive to outliers occasionally even with the $L_{1}$ error norm.
+If the model mistakenly predicts a close foreground edge pixel as a distant background point, the objective would be dominated by this incorrect prediction.
+To further improve the robustness, we apply truncation $\min(\cdot,\tau)$ to the absolute residuals.
+While this truncation introduces non-convexity, our solver remains applicable with minor adjustments.
 In summary, our alignment solver is made Robust, Optimal, and Efficient, and hence we call it the ROE solver.
-Algorithm 1 Optimal alignment solver to Eq. 4
-function solve_subproblem
-    Enumerating non-differentiable points of objective function,
-         at zero per-coordinate alignment residuals.
-    Determining extrema by derivative computation.
+Algorithm 1 Overview of ROE alignment
+function SolveSubproblem()
+    Enumerate breakpoints of the piecewise linear function.
+    Find extrema among the breakpoints by their derivatives.
     return $s^{(k)}$ with smallest objective value $l^{(k)}$ at extrema.
 end function
-for index $k=1,...,N$ do $\triangleright$ parallel computation
-    Formulate subproblem by substituting $t_{z}$ with $z^{(k)}\!-\!s\hat{z}^{(k)}$.
-    Solve scale $s^{(k)}$ and $l^{(k)}$ via SOLVE_SUBPROBLEM.
-    Obtain translation $t_{z}^{(k)}$ as $z^{(k)}\!-\!s^{(k)}\hat{z}^{(k)}$.
+for index $k=1$ to $N$ do $\triangleright$ parallel computation
+    Formulate subproblem by substituting $t_{z}$ with $z_{k}\!-s^{(k)}\hat{z}_{k}$.
+    Solve scale $s^{(k)}$ and $l^{(k)}$ via SolveSubproblem().
+    Obtain translation $t_{z}^{(k)}$ as $z_{k}\!-\!s^{(k)}\hat{z}_{k}$.
 end for
@@
 where $r_{j}$ is the radius.
-We set $r_{j}=\alpha\cdot z_{j}\cdot\frac{\sqrt{W^{2}+H^{2}}}{2\cdot f}$ where $z_{j}$ is the $z$-coordinate of $p_{j}$, $f$ is the focal length, and $W$ and $H$ are image width and height.
+We set $r_{j}=\alpha\cdot z_{j}\cdot\frac{\sqrt{W^{2}+H^{2}}}{2\cdot f}$ where $z_{j}$ is the $z$-coordinate of $p_{j}$, $f$ is the ground truth focal length, and $W$ and $H$ are image width and height.
 This way, the hyper-parameter $\alpha\in(0,1)$ approximates the proportion of the projected sphere’s diameter to the image diagonal.
-Then, we apply the aforementioned ROE alignment solver to align the point maps and compute error.
+Then, we apply the aforementioned ROE alignment solver to align the point maps with optimal $(s_{j}^{*},\mathbf{t}_{j}^{*})$ and compute error.
 For each sphere scale parameter $\alpha$, we sample a set of anchor points $\mathcal{H}_{\alpha}$ and compute the loss as
-$\mathcal{L}_{S(\alpha)}=\sum_{j\in\mathcal{H}_{\alpha}}l_{{\mathcal{S}}_{j}}=\sum_{j\in\mathcal{H}_{\alpha}}\sum_{i\in{S}_{i}}\frac{1}{z_{i}}\left\|s^{*}\hat{\mathbf{p}}_{i}+\mathbf{t}^{*}-{\mathbf{p}}_{i}\right\|_{1}$ (6)
-In practice, we use three $\alpha$ scales ranging from coarse to fine: $1/4$, $1/16$, and $1/64$, denoted as $\mathcal{L}_{S_{1}}$, $\mathcal{L}_{S_{2}}$ and $\mathcal{L}_{S_{3}}$, respectively.
+$\mathcal{L}_{S(\alpha)}=\!\sum_{j\in\mathcal{H}_{\alpha}}l_{{\mathcal{S}}_{j}}\!=\!\sum_{j\in\mathcal{H}_{\alpha}}\!\sum_{i\in{S}_{j}}\frac{1}{z_{i}}\left\|s_{j}^{*}\hat{\mathbf{p}}_{i}+\mathbf{t}_{j}^{*}-{\mathbf{p}}_{i}\right\|_{1}.$ (6)
+In practice, we use three $\alpha$ scales from coarse to fine: $1/4$, $1/16$, and $1/64$, denoted as $\mathcal{L}_{S_{1}}$, $\mathcal{L}_{S_{2}}$ and $\mathcal{L}_{S_{3}}$, respectively.
 Normal loss.
 For better surface quality, we additionally supervise the normal computed from the predicted point map with the ground truth:
-$\mathcal{L}_{N}=\sum_{i\in{\mathcal{M}}}\angle(\hat{\mathbf{n}}_{i},\mathbf{n}_{i})$ (7)
-where the normal $\mathbf{n}_{i}$ of pixel $i$ is obtained from the cross product of its adjacent edges on the image grid.
+$\mathcal{L}_{N}=\sum_{i\in{\mathcal{M}}}\angle(\hat{\mathbf{n}}_{i},\mathbf{n}_{i}),$ (7)
+where the normal $\hat{\mathbf{n}}_{i}$ of pixel $i$ is obtained from the cross product of its adjacent edges on the image grid, $\mathbf{n}_{i}$ is the ground truth, and (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ ) measures the angle difference between the two vectors.
 Mask loss.
-The distance to infinity region (e.g., sky) in outdoor scenes and plain backgrounds in object-only images are undefined.
-We add a single-channel head to our model for predicting the mask $\hat{\mathbf{M}}\in\mathbb{R}^{H\times W}$ where the distance is valid:
-$\mathcal{L}_{M}=\|\hat{\mathbf{M}}-(1-{\mathbf{M}}_{\text{inf}})\|_{2}^{2}$ (8)
+The infinity region (e.g., sky) in outdoor scenes and plain backgrounds in object-only images have undefined geometry.
+We use a single-channel output head for predicting the mask $\hat{\mathbf{M}}\in\mathbb{R}^{H\times W}$ of valid points:
+$\mathcal{L}_{M}=\|\hat{\mathbf{M}}-(1-{\mathbf{M}}_{\text{inf}})\|_{2}^{2},$ (8)
 where ${\mathbf{M}}_{\text{inf}}$ is the infinity mask label.
 For synthetic data, the ground truth masks are readily available, while for real outdoor scenes, we use SegFormer [64] to obtain sky masks.
-At inference time, we use a threshold of $0.5$ to binarize the mask prediction. Recovery camera focal and shift involves masked points only, excluding the regions of meaningless geometry.
+At inference time, the predicted mask is binarized with a threshold of $0.5$.
+The recovery of focal and shift involves masked points, excluding the regions of invalid geometry.
 3.3 Training data
-Recently, numerous synthetic and real RGBD datasets have been utilized in large-scale training for monocular geometry estimation.
-In this work, we collected $21$ datasets from various domains to train our model.
-The total number of collected frames is around $9$ million.
-Details of these datasets and their mixing ratios can be found in the suppl.
+We collected 21 publicly available datasets from various domains [21, 62, 3, 14, 69, 35, 74, 51, 57, 45, 58, 40, 34, 19, 27, 37, 76, 47, 60, 25, 12], totaling around 9 million frames for training our model.
+Since the quality of data labels are varied, we apply tailored loss function combinations for different data sources.
+Higher-quality and less noisy data are assigned with finer-level loss functions.
+To address imperfections and potential outliers in real labels, we exclude the top $5\%$ of per-pixel loss values for real data.
+Following DINOv2 [41], we balance the weights of collected datasets via image retrieval from a curated image set.
+Details of training datasets can be found in the suppl.
 materials.
-Since the quality of data labels are varied, we apply different loss function combinations tailored for different data sources.
-As shown in Table 1, higher quality and less noisy data are assigned with finer-level loss functions.
-To address imperfections and potential outliers in real labels, we exclude the top $5\%$ of per-pixel loss values for real data.
-Following DINOv2 [40], we rebalance the collected datasets using image retrieval from a curated image set to address uneven or biased data distributions.
-See the suppl.
-materials for details.
-Table 1: Label quality and applied losses for different training sources.
-Source Label quality Applied losses
-Accuracy Range Density $\mathcal{L}_{G}$ $\mathcal{L}_{S_{1}}$ $\mathcal{L}_{S_{2}}$ $\mathcal{L}_{S_{3}}$ $\mathcal{L}_{N}$ $\mathcal{L}_{M}$
-Synthetic Perfect $\infty$ Dense
-SfM/MV Recon High $\infty$ Dense&Partial
-LiDAR/Laser High $\sim 100$m Sparse
-Kinect Medium $\sim 10$m Dense
 4 Experiments
@@
 […]
-Scale-invariant depth
+     Scale-invariant depth map
 LeReS 12.1 82.5 19.3 64.3 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.1 18.1 69.9 18.5 69.5 7.31
@@
 Ours 3.48 98.4 4.27 97.8 3.38 98.9 3.45 97.1 1.49 100 19.4 73.4 9.11 90.6 4.85 94.8 6.17 93.9 1.62
-Affine-invariant depth
+     Affine-invariant depth map
 Marigold 4.63 97.3 7.29 93.8 6.08 96.3 4.35 97.2 2.78 99.9 21.2 75.0 14.6 80.5 6.34 94.3 8.41 91.8 2.25
 GeoWizard 4.69 97.4 8.14 92.5 6.90 94.0 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.44 90.7 2.69
-Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.56
-Affine-invariant disparity
+Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.00
+     Affine-invariant disparity map
 MiDaS V3.1 4.62 98.1 6.30 94.8 5.80 96.8 4.75 97.4 1.88 100 21.4 73.0 14.4 82.7 6.06 94.9 8.15 92.2 3.69
@@
 DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 2.50
-Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.944 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.06
-Table 3: Quantitative results for depth map estimation. Gray numbers denote models trained on respective benchmarks and thus excluded from ranking.
+Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.06
+Table 2: Quantitative results for depth map estimation.
+Gray numbers denote models trained on respective benchmarks and thus excluded from ranking.
+An extensive comparison is found in the suppl.
+materials.
 Method NYUv2 ETH3D iBims-1 Average
@@
 Med. affine-inv. 11.5 88.5 9.36 89.7 10.2 90.9 9.11 89.7 7.38 92.4 8.53 92.0
-Full w/o clip 10.1 90.4 8.13 91.4 6.94 94.0 7.94 91.4 6.31 93.5 7.45 93.1
+Full w/o truncation 10.1 90.4 8.13 91.4 6.94 94.0 7.94 91.4 6.31 93.5 7.45 93.1
 Full w/o $\mathcal{L}_{S}$ 10.1 90.2 8.10 91.8 7.51 93.5 7.94 91.2 6.31 93.6 7.47 93.2
@@
 Implementation details.
-We employ a ViT [15] encoder and a light-weight CNN-based upsampler as the decoder.
-The encoder is pretrained with DINOv2 [40].
-The initial learning rates for the encoder and decoder are set to $5\times 10^{-6}$ and $5\times 10^{-5}$, with a decay factor of $5$ applied for every 100K iterations.
+We use a ViT [15] encoder pretrained with DINOv2 [41] and a lightweight CNN-based upsampler as the decoder.
+The encoder and decoder have initial learning rates of $5\times 10^{-6}$ and $5\times 10^{-5}$, with a decay factor of 5 every 100K iterations.
 The training batch size is $256$.
-Our model is trained on images with varying aspect ratio and resolutions, where aspect ratios are uniformly sampled between $1:2$ and $2:1$, and pixel counts range from 250K to 500K.
+Our model is trained on images with aspect ratios uniformly sampled between $1:2$ and $2:1$, and pixel counts from 250K to 500K.
 We apply several image augmentations, including color jittering, Gaussian blurring, JPEG compression-decompression, and random cropping.
@@
 4.1 Quantitative Evaluations
-We assess the zero-shot performance of our model and compare it with several state-of-the-art methods for monocular point map estimation [69, 41, 57] and depth map estimation [65, 66, 6, 26, 5, 20, 29].
-To ensure a fair comparison, all methods utilize ViT-Large [15] as backbone, except for LeReS (which employs ResNeXt101 [63]) as well as GeoWizard and Marigold (both use StableDiffusion V1 [45]).
-For a valid evaluation, the scale and shift of the predicted point maps and depth maps are first aligned with the ground truth when necessary.
+We assess the zero-shot performance of our model and compare it to several state-of-the-art methods for monocular point map estimation [71, 42, 59] and depth map estimation [67, 68, 6, 26, 5, 20, 29].
+To ensure a fair comparison, all methods utilize ViT-Large [15] as backbone, except for LeReS (using ResNeXt101 [65]) as well as GeoWizard and Marigold (using StableDiffusion V1 [46]).
+The scale and shift of the predicted point maps and depth maps are aligned with the ground truth when necessary.
 We also evaluate the FOV prediction results against several recent methods [28, 77].
-As detailed below, our method achieves superior results across all evaluation metrics for various tasks on different benchmarks.
+As detailed below, our method achieves superior results across various tasks and benchmarks.
 Point map estimation.
@@
 material for more preprocessing details).
-For evaluation metrics, we use the relative point error Relp, i.e., $\|\mathbf{p}-\hat{\mathbf{p}}\|_{2}/\|\hat{\mathbf{p}}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\mathbf{p}-\hat{\mathbf{p}}\|_{2}/\|\hat{\mathbf{p}}\|_{2}<0.25$.
-We evaluate our method and compare it with LeReS [69], UniDepth [41], and DUSt3R [57].
-LeReS and UniDepth predict depth and camera parameters from single images.
-DUSt3R is designed for two-view input and directly outputs camera-space point maps.
-It accepts two identical images for handling monocular scenarios.
-We evaluate both scale-invariant point maps in camera space and affine-invariant point maps.
-Table 2 demonstrates that our method outperforms existing methods across most benchmarks, achieving the lowest average Relp for both point maps representations.
-To further evaluate local geometry accuracy, we assess affine-invariant point maps within local regions defined by object segmentation masks, either from the dataset labels or obtained using Segment Anything [30].
-Since per-region point map evaluation demands high-quality ground truth labels as well as the composition of multiple objects in a scene, NYUv2, KITTI, and GSO are excluded.
-Compared to the previous methods, our method significantly reduces the region-wise Relp by a large margin: from $7.96$ to $5.33$, achieving approximately a $30\%$ error reduction.
+For evaluation metrics, we use the relative point error Relp, i.e., $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}<0.25$.
+We compare our method with LeReS [71], UniDepth [42], and DUSt3R [59], evaluating both camera-space scale-invariant point maps and affine-invariant point maps.
+Table 1 demonstrates that our method outperforms existing methods, achieving the lowest average Relp and highest $\delta_{1}^{\text{p}}$ for both point map representations.
+To further assess local geometry accuracy, we evaluate affine-invariant point maps within regions defined by object segmentation masks, sourced from dataset labels or Segment Anything [30].
+Due to the need for high-quality ground truth and scenes with multiple objects, NYUv2, KITTI, and GSO are excluded.
+Our method achieves a substantial reduction in region-wise Relp, lowering the error from $7.96$ to $5.33$—an approximately $30\%$ improvement over previous methods.
 Depth map estimation.
@@
 We adopt absolute relative error Reld, i.e., $|z-\hat{z}|/z$, and percentage of inliers $\delta_{1}^{d}$ with $\max(z/\hat{z},\hat{z}/z)<1.25$ as the evaluation metrics.
-To conduct a comprehensive comparison with different monocular depth estimation methods, we evaluate scale-invariant depth [69, 5, 57, 26, 41], affine-invariant depth [29, 20], and affine-invariant disparity [6, 65, 66].
-The $z$ coordinates in our predicted point maps represent affine-invariant depth.
-By utilizing the camera shifts derived from Eq.
-2, we convert these coordinates into scale-invariant depth.
+For a comprehensive comparison with different monocular depth estimation methods, we evaluate scale-invariant depth [71, 5, 59, 26, 42], affine-invariant depth [29, 20], and affine-invariant disparity [6, 67, 68].
+The $z$ coordinates in our predicted point maps represent affine-invariant depth, which can be converted to scale-invariant depth using the shift derived from Eq.
+2.
 For affine-invariant disparity, we take the inverse of our scale-invariant depth predictions.
-As shown in Table 3, our method achieves the best performance across most benchmarks and exhibits the lowest average Reld for all depth representations.
+As shown in Table 2, our method exhibits the lowest average Reld for all depth representations.
 Since existing methods with scale-invariant depth representations can be converted to affine-invariant depth and disparity, we provide an extensive comparison in the suppl.
@@
 To evaluate the accuracy of the estimated camera FOV, we select benchmarks with vertical $\text{vertical FOV}\geq 45^{\circ}$, including NYUv2 [39], ETH3D [50] and iBims-1 [31], and augment the FOV by randomly center cropping 50$\sim$100from the original images.
-All input images are undistorted and have a centered principal point.
-We compare our model to point map estimation methods [69, 41, 57] as well as two learning-based camera calibration methods: Perspective [28] and WildCam [75].
-For fair assessments of zero-shot performance, we utilize their pre-trained checkpoints on in-the-wild datasets.
-Mean and median FoV errors are reported in Table 5.
-Our method achieves a mean error of $2.91^{\circ}$ and a median error of $2.21^{\circ}$ on average of the three FoV benchmarks, outperforming both camera calibration methods and point map estimation methods.
+All input images are undistorted with a centered principal point.
+We compare our model to point map estimation methods [71, 42, 59] as well as two learning-based camera calibration methods: Perspective Fields [28] and Wild Camera [77].
+For fair zero-shot evaluation, we utilize their pre-trained checkpoints on in-the-wild datasets.
+Table 4 reports mean and median FOV errors.
+Our method achieves a mean error of $2.91^{\circ}$ and a median error of $2.21^{\circ}$ on average of the three FOV benchmarks, outperforming both camera calibration and point map estimation methods.
 4.2 Qualitative comparison
@@
 7.
-For depth prediction methods that do not recover camera parameters, we lift their results to 3D points using ground-truth intrinsics when available.
-Upon visual inspection, our method produces significantly less-distorted point maps compared to others, demonstrating its superior generalizabiltiy and accuracy.
-For example, as shown in the fourth row of Fig.
-7, other methods result in noticeable stretching and deformation of the car shape.
-Although [66] produces visually sharp disparity maps, distorted 3D geometry can be observed even when provided with ground-truth camera intrinsics.
+For depth prediction methods without camera parameters estimation, we lift their results to 3D points using ground-truth intrinsics when available.
+Upon visual inspection, our method produces significantly less-distorted point maps compared to others, demonstrating its superior generalizability and accuracy.
+For example, as shown in the third row of Fig.
+7, other methods show noticeable stretching and deformation of the car shape.
+Although [68] produces visually sharp disparity maps, distorted 3D geometry is evident even with ground-truth camera intrinsics.
 In contrast, our method yields a more regular and well-formed car structure.
+More comparisons can be found in the suppl.
+materials.
 Figure 7: Qualitative comparison of point and disparity maps.
-*: for methods without camera intrinsics prediction, we use ground-truth camera intrinsics or disparity shifts to lift their results into 3D points when available.
-The first four cases are samples from our test datasets, while the last two are random “in-the-wild” images.
+The first three cases are from our test datasets, while the last two rows are random “in-the-wild” images.
+*: for methods without camera intrinsics prediction, we use ground-truth camera intrinsics (and disparity shifts) to lift their results into 3D points when available, or use our intrinsics prediction for the “in-the-wild” images.
+Best view with zoom.
 4.3 Ablation Study
-We conducted comprehensive ablation studies on our point map representation, alignment strategy and loss function designs. For efficiency, we use the ViT-base model as backbone for these experiments.
+We conducted comprehensive ablation studies on our point map representation, alignment strategy and loss function designs. ViT-base backbones are used here for efficiency.
 Point map representation.
-Our method predicts affine-invariant point maps.
-Here we compare this representation with two variants, one scale-invariant point map in camera space, and another that predicts a depth maps and camera parameters.
-For the former, we apply the proposed ROE alignment solver to solve the scale parameter only.
+We compare our affine-invariant point map representation with two variants, one scale-invariant point map in camera space, and another that predicts a depth map and camera parameters.
+For the former, we apply the proposed ROE alignment to solve the scale parameter only.
 For the latter, we adopt the SI-log loss [18] for the depth map and the MSE loss for the ray direction map, similar to UniDepth [42].
-The first three rows of Table 5 show the results of different approaches.
-In these experiments, only the global ground-truth supervision is applied during training.
-Our proposed affine-invariant representation outperforms the other variants across various tasks.
-This can be attributed to the fact that the other two variants suffer from focal-distance ambiguity issue during training which jeopardizes their performance.
-Optimal alignment.
-To evaluate the effectiveness of our optimal alignment, we compare it to a previously-used simple alignment method [42], i.e., we normalize the predicted and ground-truth point maps by their median $z$: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})$ equals to $(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})$ is computed as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$.
+Results are shown in the first three rows of Table 4, where only the global ground-truth supervision is applied in these experiments.
+Our proposed affine-invariant representation outperforms the others across various tasks.
+This is attributed to the fact that the other two variants suffer from focal-distance ambiguity issue during training, which hinders their performance.
+ROE alignment.
+To evaluate the effectiveness of our optimal alignment, we compare it to a previously-used simple alignment method [43], i.e., the predicted and ground-truth point maps are normalized by their median: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})$ equals to $(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})$ is computed as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$.
 As shown in the third and fourth rows of Table 4, our proposed alignment method consistently outperforms the median alignment strategy across all tasks by large margins, demonstrating the critical role of our proposed optimal alignment.
+We ablate the truncation of residuals in alignment objects, which degrades performance (see the fifth row in Table 4). In practice, adding truncation also improves training stability.
 Multi-scale local geometry loss.
@@
 Infinity mask.
-Our method predicts a valid region mask to handle regions at infinity.
+Our method predicts a valid region mask to handle infinity regions.
 As shown in Fig.
-6 (b), our method correctly predicts the sky regions, for which the predicted point values will be erroneous if the masks are removed.
-We further compare our mask prediction design with another strategy for addressing infinity.
-Specifically, we assign a large value, i.e., $1,000$ times the average distance of valid regions, to the infinity regions.
+6 (b), it correctly identifies sky regions, where the output point values would be erroneous if the masks were removed.
+Another strategy for addressing infinity is to assign a large value, i.e., 1,000×1,000$\times$1 , 000 × the average distance, to these regions.
 Fig.
-6 (b) shows that this strategy compromises the accuracy of foreground predictions as the network struggles to generate large values for infinity.
-Incorporating a separate mask prediction offers a convenient solution for masking out infinity and retaining accurate foreground geometry learning.
+6 (b) shows that this strategy compromises the foreground accuracy as the network struggles with large values for infinity.
+Incorporating a separate mask prediction effectively handles infinity while retaining accurate foreground geometry.
 5 Conclusion
@@
 Our key insight is to design effective supervision which is largely neglected by previous methods.
-We propose a new affine-invariant point map representation which precludes ambiguous supervision during training.
-A robust, optimal, and efficient point map alignment solver is introduced for accurate global shape learning, together with a multi-scale local geometry loss designed for precise local geometry supervision.
-Trained on a large dataset collection, our model demonstrates strong generalizability and high accuracy for open-domain images.
-It significantly outperforms previous methods across various tasks and benchmarks.
-We believe our method marks a significant advancement in monocular 3D geometry estimation and can serve as a robust foundation model for various vision applications.
+We introduce an affine-invariant point map representation which precludes ambiguous supervision during training.
+A robust, optimal, and efficient point map alignment solver is introduced for effective shape learning, along with a multi-scale geometry loss for precise local supervision.
+Trained on a large dataset collection, our model demonstrates strong generalizability and significantly outperforms previous methods across various tasks and benchmarks.
+We believe our method marks a significant advancement in monocular 3D geometry estimation and can serve as a robust foundational model for various applications.
 Supplementary Material
@@
 The focal length and shift is obtained by minimizing the projection error,
-$\min_{f,t_{z}}\sum_{i\in\mathcal{M}}\left({fx_{i}\over z_{i}+t_{z}^{\prime}}-u_{i}\right)^{2}+\left({fy_{i}\over z_{i}+t_{z}^{\prime}}-v_{i}\right)^{2}$ (9)
-which can be further reduced to have a single variable $t_{z}^{\prime}$ by substituting $f$ with its close-form solution with respect to $t_{z}^{\prime}$
-$f={\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}\right)u_{i}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}\right)v_{i}\over\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}\right)^{2}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}\right)^{2}}$ (10)
+$\min_{f,t_{z}}\sum_{i\in\mathcal{M}}\left({fx_{i}\over z_{i}+t_{z}^{\prime}}-u_{i}\right)^{2}+\left({fy_{i}\over z_{i}+t_{z}^{\prime}}-v_{i}\right)^{2},$ (9)
+which can be further reduced to have a single variable $t_{z}^{\prime}$ by substituting $f$ with its close-form solution with respect to $t_{z}^{\prime}$,
+$f={\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}^{\prime}\right)u_{i}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}^{\prime}\right)v_{i}\over\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}^{\prime}\right)^{2}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}^{\prime}\right)^{2}}.$ (10)
 We use a numerical solver for this least squares problem with Levenberg-Marquardt algorithm [38] implemented by SciPy [55] package.
@@
 In our practice, it typically converges within 10 iterations in around 3ms.
-A.2 Optimal alignment
+A.2 ROE Alignment
 We will first introduce an algorithm to a simpler subproblem then derive the solution to either with or without the constraint of $t_{x}=t_{y}=0$ (1D-shift case or 3D-shift case, respectively).
-Subproblem (w/o clipping)
-Consider the optimization objective with respect to scale $s$ only, denoted as $l_{0}(s)$. We omit the mask $\mathcal{M}$ for simplicity and denote $N$ as the number of valid points.
-$\min_{s}l_{0}(s)=\min_{s}\sum_{i=1}^{N}w_{i}|s\hat{x}_{i}-x_{i}|$ (11)
+Subproblem (w/o truncation).
+Consider the optimization objective with respect to scale $s$ only, denoted as $l_{0}(s)$. We omit the mask $\mathcal{M}$ for simplicity and denote $N$ as the number of valid points:
+$\min_{s}l_{0}(s)=\min_{s}\sum_{i=1}^{N}w_{i}|s\hat{x}_{i}-x_{i}|,$ (11)
 where $w_{i}>0$ and $\hat{x}_{i}>0$ without loss of generality.
@@
 The minimum occurs where its left-hand $l_{0}^{\prime-}(s)$ derivative and right-hand derivative $l_{0}^{\prime+}(s)$ have opposite signs or one of them is zero,
-$\begin{split}l_{0}^{\prime-}(s)&=\sum_{{x_{i}\over\hat{x}_{i}}<s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\sum_{i=1}^{N}w_{i}\hat{x}_{i}\\ l_{0}^{\prime+}(s)&=\sum_{{x_{i}\over\hat{x}_{i}}\leq s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\sum_{i=1}^{N}w_{i}\hat{x}_{i}\end{split}$ (12)
-$l_{0}^{\prime-}(s)$ and $l_{0}^{\prime+}(s)$ differ at $\left\{x_{i}\over\hat{x}_{i}\right\}$.
-One might first sort $\left\{x_{i}\over\hat{x}_{i}\right\}$ and compute the prefix summations of $\{w_{i}\hat{x}_{i}\}$.
-Then the derivatives can be evaluated in $O(1)$ time for each point in $\left\{x_{i}\over\hat{x}_{i}\right\}$.
-Finally, $\hat{x}_{i}\over x_{i}$ such hat ł$\l_{0}^{\prime-}\left(x_{i}\over\hat{x}_{i}\right)<0\leq\l_{0}^{\prime+}\left(x_{i}\over\hat{x}_{i}\right)$ is a deterministic minimum point.
-Subproblem (w/ clipping)
-We clip each residual term to suppress the contribution of outliers. Specifically, the clipped objective is
-$\min_{s}l_{1}(s)=\min_{s}\sum_{i=1}^{N}\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ (13)
-For each item $l_{1,i}(s)=\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ in the equation, the one-sided derivatives are
-$\begin{split}l_{1,i}^{\prime-}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.\\ l_{1,i}^{\prime+}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.\end{split}$ (14)
+$\begin{split}l_{0}^{\prime-}(s)&=\!\!\sum_{{x_{i}\over\hat{x}_{i}}<s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\!\!\sum_{i=1}^{N}w_{i}\hat{x}_{i},\\ l_{0}^{\prime+}(s)&=\!\!\sum_{{x_{i}\over\hat{x}_{i}}\leq s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\!\!\sum_{i=1}^{N}w_{i}\hat{x}_{i}.\end{split}$ (12)
+$l_{0}^{\prime-}(s)$ and $l_{0}^{\prime+}(s)$ differ at $\{{x_{i}\over\hat{x}_{i}}\}$.
+First, we sort $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$ and compute the prefix summations of $\{w_{i}\hat{x}_{i}\}_{i=1}^{N}$.
+This allows us to evaluate the derivatives in $O(1)$ time for each point in $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$.
+Finally, $\hat{x}_{i}\over x_{i}$ such that ł$\l_{0}^{\prime-}({x_{i}\over\hat{x}_{i}})\leq 0\leq\l_{0}^{\prime+}({x_{i}\over\hat{x}_{i}})$ is the minimum point.
+The solution is outlined in Algorithm 2.
+Algorithm 2 ROE alignment subproblem w/o truncation
+input: arrays $\hat{X}[1..n]$, $X[1..n]$, $W[1..n]$
+output: optimal scale $s^{*}$ and objective value $l^{*}$ to Eq. 11
+function SolveSubproblem($\hat{X}$, $X$, $W$)
+       sort arrays $\hat{X}$, $X$, $W$ by $X[i]/\hat{X}[i]$
+       $Q[1..n]$leftarrow accumulated sum of $W*\hat{X}$
+       $D[0..n]$leftarrow $\{-Q[n]\}\cup\{2\cdot Q[i]-Q[n]\}_{i=1}^{n}$
+       $i^{*}$leftarrow the first $i$ s.t. $D[i-1]\leq 0\leq D[i]$
+       $s^{*}$leftarrow $X[i^{*}]/\hat{X}[i^{*}]$
+       $l^{*}$leftarrow objective function value at $s^{*}$.
+       return $s^{*}$, $l^{*}$.
+end function
+Subproblem (w/ truncation).
+We truncate each residual term to suppress outliers. The truncated objective is
+$\min_{s}l_{1}(s)=\min_{s}\sum_{i=1}^{N}\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|),$ (13)
+where $\tau$ is set to $1$ in all our experiments. For each item $l_{1,i}(s)=\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ in the equation, the one-sided derivatives are
+$\begin{split}l_{1,i}^{\prime-}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.,\\ l_{1,i}^{\prime+}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right..\end{split}$ (14)
 Therefore, the one-sided derivatives of $l_{0}(s)$ are
-$\begin{split}l_{1}^{\prime-}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime-}(s)\\ &=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}\end{split}$ (15)
-$\begin{split}l_{1}^{\prime+}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime+}(s)\\ &=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}\end{split}$ (16)
-One can evaluate $l_{1}^{\prime}(s)$ efficiently through first binary searching on the sorted arrays $\left\{x_{i}\over\hat{x}_{i}\right\}$, $\left\{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}\right\}$ and $\left\{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}\right\}$ and then indexing the precomputed prefix summations of $\{w_{i}\hat{x}_{i}\}$ in the associated orders.
+$\begin{split}l_{1}^{\prime-}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime-}(s)=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i},\end{split}$ (15)
+$\begin{split}l_{1}^{\prime+}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime+}(s)=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}.\end{split}$ (16)
+Lemma 1 shows that the minimum of the objective function in Eq.
+13 is still achieved at one of the points in the set $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$, despite the function is non-convex and may contain local minima.
+Solving the subproblem requires two steps, as outlined in Algorithm 3.
+The first step is to identify all extrema in $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$ that satisfy $l_{0}^{\prime-}({x_{i}\over\hat{x}_{i}})<0\leq l_{0}^{\prime+}({x_{i}\over\hat{x}_{i}})$ by evaluating the derivative values.
+This can be done efficiently through first binary searching on the sorted arrays $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$, $\{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\}_{i=1}^{N}$ and $\{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}\}_{i=1}^{N}$ and then indexing the prefix summations of $\{w_{i}\hat{x}_{i}\}$ in the associated orders.
+This step has a complexity of $O(N\log N)$.
+The second step involves computing the objective values at these extrema and determining the minimum, which takes $O(Nn_{e})$ time, where $n_{e}$ is the number of extrema.
+As $n_{e}$ approximates the number of outliers, it is typically a small constant in practice.
 Lemma 1.
-There exists at least one pair of $(k^{*},s^{*})$ such that $s^{*}x_{k^{*}}-\hat{x}_{k^{*}}=0$ and $s^{*}$ minimizes Eq. 13
+There exists at least one pair of $(k^{*},s^{*})$ such that $s^{*}\hat{x}_{k^{*}}-x_{k^{*}}=0$ and $s^{*}$ minimizes Eq. 13.
 Proof.
@@
 Given $l_{0}^{\prime+}(s^{*})>l_{0}^{\prime-}(s^{*})$, there exists an index $k^{*}$ such that $s^{*}=\hat{x}_{k^{*}}/x_{k^{*}}$, because
-$\begin{split}0<&l_{1}^{\prime+}(s^{*})-l_{1}^{\prime-}(s^{*})\\ =&2\sum_{{\hat{x}_{i}\over x_{i}}=s^{*}}w_{i}x_{i}-\!\!\!\!\!\!\sum_{{w_{i}\hat{x}_{i}-\tau\over w_{i}x_{i}}=s^{*}}\!\!\!\!w_{i}x_{i}-\!\!\!\!\!\!\sum_{{w_{i}\hat{x}_{i}+\tau\over w_{i}x_{i}}=s^{*}}\!\!\!\!w_{i}x_{i}\\ <&2\sum_{{\hat{x}_{i}\over x_{i}}=s^{*}}w_{i}x_{i}\end{split}$ (17)
-Lemma 1 has proved that the minimum point of the objective function in Eq.
-13 still occurs among $\left\{x_{i}\over\hat{x}_{i}\right\}$, although the function is non-convex and may contain local minima.
-Therefore, it takes two steps to solve the subproblem.
-First, find all extrema among $\left\{x_{i}\over\hat{x}_{i}\right\}$ such that $l_{0}^{\prime-}\left(x_{i}\over\hat{x}_{i}\right)<0\leq l_{0}^{\prime+}\left(x_{i}\over\hat{x}_{i}\right)$ by evaluating the derivative values, which takes $O(N\log N)$ time.
-Second, evaluate the objective values of those extrema and then find the minimum, which takes $O(Nn_{e})$ time.
-$n_{e}$ is the number of extrema, approximating the number of outliers , which is expected to be a small constant in practice.
-Alignment with 1D shift
-Recall the objective in Eq. 4. Let $w_{i}$ be $1/z_{i}$. We apply clipping to each absolute term and rewrite it as follows
-$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-x_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|)\end{split}$ (18)
+$\begin{split}0<&l_{1}^{\prime+}(s^{*})-l_{1}^{\prime-}(s^{*})\\ =&2\sum_{{x_{i}\over\hat{x}_{i}}=s^{*}}w_{i}\hat{x}_{i}-\!\!\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}=s^{*}}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}=s^{*}}\!\!\!\!w_{i}\hat{x}_{i}\\ \leq&2\sum_{{x_{i}\over\hat{x}_{i}}=s^{*}}w_{i}\hat{x}_{i}.\end{split}$ (17)
+Algorithm 3 ROE alignment subproblem w/ truncation
+input: arrays $\hat{X}[1..n]$, $X[1..n]$, $W[1..n]$, float $\tau$
+output: the optimal scale $s^{*}$, objective value $l^{*}$ to Eq. 13
+function SolveSubproblem($\hat{X}$, $X$, $W$, $\tau$)
+       $A[1..n]$leftarrow $X/\hat{X}$
+       $B[1..n]$leftarrow $(W*X-\tau)/(W*\hat{X})$
+       $C[1..n]$leftarrow $(W*X+\tau)/(W*\hat{X})$
+       for each array $\mathcal{A}$ in $\{A,B,C\}$ do
+           sort $\mathcal{A}$ and obtain sorted indices $I_{\mathcal{A}}[1..n]$
+           $Q_{\mathcal{A}}[1..n]$leftarrow accumulated sum of $\{W\hat{X}[I_{\mathcal{A}}[i]]\}_{i=1}^{n}$
+       end for
+       Initialize $I_{E}$ as empty set
+       for $i=1$ to $n$ do $\triangleright$ parallel computation
+           for each array $\mathcal{A}$ in $\{A,B,C\}$ do
+               $j_{\mathcal{A}}^{-}$leftarrow the last $j$ s.t. $\mathcal{A}[j]<X[i]/\hat{X}[i]$
+               $j_{\mathcal{A}}^{+}$leftarrow the last $j$ s.t. $\mathcal{A}[j]\leq X[i]/\hat{X}[i]$
+           end for
+           $d^{-}$leftarrow $2\cdot Q_{A}[j_{A}^{-}]-Q_{B}[j_{B}^{-}]-Q_{C}[j_{C}^{-}]$
+           $d^{+}$leftarrow $2\cdot Q_{A}[j_{A}^{+}]-Q_{B}[j_{B}^{+}]-Q_{C}[j_{C}^{+}]$
+           if $d^{-}<0\leq d^{+}$ then append $i$ to $I_{E}$
+       end for
+       Initialize $l[1..N]$ with $\infty$
+       for $i$ in $I_{E}$ do $\triangleright$ parallel computation
+           $s$leftarrow $X[i]/\hat{X}[i]$
+           $l[i]$leftarrow objective function value at $s$
+       end for
+       $i^{*}$leftarrow index of the minimum in $l[i],i\in I_{E}$
+       $s^{*}$leftarrow $X[i^{*}]/\hat{X}[i^{*}]$, $l^{*}$leftarrow $l[i^{*}]$
+       return $s^{*}$, $l^{*}$
+end function
+Alignment with 1D shift.
+Recall the alignment objective and let $w_{i}$ be $1/z_{i}$. We rewrite it as follows:
+$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}[w_{i}|s\hat{x}_{i}\!-x_{i}|+w_{i}|s\hat{y}_{i}\!-\!y_{i}|+w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|],\end{split}$ (18)
+or apply truncation to each absolute residual term
+$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}[\min(\tau,w_{i}|s\hat{x}_{i}\!-x_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|)].\end{split}$ (19)
+The proposed solution is outlined in Algorithm 4, with proof as follows. The corresponding subproblem solver is selected based on whether truncation is applied.
 Lemma 2.
@@
 Denote the objective as $l_{2}(s,t_{z})$,
-$\begin{split}l_{2}(s,t_{z})=&\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{x}_{i}-x_{i}|}\right)\\ &+\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{y}_{i}-y_{i}|}\right)\\ &+\sum_{i=1}^{n}\min\left(\tau,w_{i}{|t_{z}-(z_{i}-s\hat{z}_{i})|}\right)\end{split}$ (19)
-Given arbitrary $s$, using Lemma 1, there exists at least one pair $(t_{z},k)$ such that $t_{z}-(z_{k}-s\hat{z}_{k})=0$ and $t_{z}$ minimizes $\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{z}_{i}+t_{z}-z_{i}|}\right)$, hence minimizes $l(s,t_{z})$ as the rest parts are constant with regard to $t_{z}$.
+$\begin{split}l_{2}(s,t_{z})&=\sum_{i=1}^{n}[\min\left(\tau,w_{i}{|s\hat{x}_{i}-x_{i}|}\right)+\min\left(\tau,w_{i}{|s\hat{y}_{i}-y_{i}|}\right)]\\ &+\sum_{i=1}^{N}\min\left(\tau,w_{i}{|t_{z}-(z_{i}-s\hat{z}_{i})|}\right)].\end{split}$ (20)
+Given arbitrary $s$, using Lemma 1, there exists at least one pair $(t_{z},k)$ such that $t_{z}-(z_{k}-s\hat{z}_{k})=0$ and $t_{z}$ minimizes $\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{z}_{i}+t_{z}-z_{i}|}\right)$, hence minimizes $l_{2}(s,t_{z})$ as the rest parts are constant with regard to $t_{z}$.
 Therefore, a solution $s^{*}$ is always associated with corresponding $(t^{*}_{z},k^{*})$ such that that $s^{*}\hat{z}_{k^{*}}+t^{*}_{z}-z_{k^{*}}=0$.
 ∎
-Lemma 2 allows us to reduce Eq. 18 to the subproblem with respect to some index $k$. For each possible index $k$, the objective is formed as
-$\begin{split}\min_{s}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-\!sx_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s(\hat{z}_{i}-\hat{z}_{k})-(z_{i}-z_{k})|)\end{split}$ (20)
+Lemma 2 allows us to reduce Eq. 19 to the subproblem with respect to some index $k$. For each possible index $k$, the objective is formed as:
+$\begin{split}\min_{s}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-\!sx_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s(\hat{z}_{i}-\hat{z}_{k})-(z_{i}-z_{k})|),\end{split}$ (21)
 which is solvable in $O(N\log N)$ complexity. We enumerate all possible indices for $k$ and find the minimum. Therefore, the total time complexity is $O(N^{2}\log N)$.
 In our implementation, the point map is resized to low resolution ($64\times 64$) for alignment, with $N=4096$ at most. The algorithm is further parallelized with tensor operations on GPUs.
-Alignment with 3D shift
+Algorithm 4 ROE alignment w/ 1-D shift
+input: point arrays $\hat{P}[1..N,1..3]$, $P[1..N,1..3]$,
+   weight array $W[1..N]$
+output: the optimal scale $s^{*}$, shift $t_{z}^{*}$,
+   objective value $l^{*}$ to Eq. 18 or Eq. 19
+$W[1..3N]$leftarrow repeat each element in $W$ 3 times
+Initialize arrays $s[1..N]$, $l[1..N]$, $t_{z}[1..N]$
+for $k=1$ to $N$ do $\triangleright$ parallel computation
+[…]
+[…]
+       $(s[k],l[k])$leftarrow SolveSubproblem($\hat{X}$,$X$,$W$)
+       $t_{z}[k]$leftarrow $P[k,3]-s[k]\cdot\hat{P}[k,3]$
+end for
+$k^{*}$leftarrow index of the minimum in $l[1..N]$
+$s^{*}$leftarrow $s[k^{*}]$, $l^{*}$leftarrow $l[k^{*}]$, $t_{z}^{*}$leftarrow $t_{z}[k^{*}]$
+return $s^{*}$, $t_{z}^{*}$, $l^{*}$
+Alignment with 3D shift.
+We apply truncation and rewrite the objective as follows:
+$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}&\min(\tau,w_{i}|s\hat{x}_{i}\!+t_{x}-x_{i}|)\\ &+\min(\tau,w_{i}|s\hat{y}_{i}+t_{y}-y_{i}|)\\ &+\min(\tau,w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|).\end{split}$ (22)
 Similarly to the proof of Lemma 2, there exists at least one group $(k_{1}^{*},k_{2}^{*},k_{3}^{*},s^{*},\mathbf{t}^{*})$ such that $s^{*}\hat{x}_{k_{1}^{*}}+t_{x}^{*}-x_{k_{1}^{*}}=0$, $s^{*}\hat{y}_{k_{2}^{*}}+t_{y}^{*}-y_{k_{2}^{*}}=0$, $s^{*}\hat{z}_{k_{3}^{*}}+t_{z}^{*}-z_{k_{3}^{*}}=0$, and $(s^{*},\mathbf{t}^{*})$ minimizes the objective.
@@
 The effectiveness of the approximated solution has been empirically validated.
-Appendix B Data details
+Algorithm 5 ROE alignment w/ 3-D shift
+input: point arrays $\hat{P}[1..N,1..3]$, $P[1..N,1..3]$,
+   weight array $W[1..N]$
+output: the optimal scale $s^{*}$, shift $\mathbf{t}^{*}$,
+   objective value $l^{*}$ to Eq. 22
+$W[1..3N]$leftarrow repeat each element in $W$ 3 times
+Initialize arrays $s[1..N]$, $l[1..N]$, $\mathbf{t}[1..N,3]$
+for $k=1$ to $N$ do $\triangleright$ parallel computation
+[…]
+[…]
+       $(s[k],l[k])$leftarrow SolveSubproblem($\hat{X}$,$X$,$W$)
+       $\mathbf{t}[k]$leftarrow $P[k]-s[k]\cdot\hat{P}[k]$
+end for
+$k^{*}$leftarrow index of the minimum in $l[1..N]$
+$s^{*}$leftarrow $s[k^{*}]$, $l^{*}$leftarrow $l[k^{*}]$, $\mathbf{t}^{*}$leftarrow $\mathbf{t}[k^{*}]$
+return $s^{*}$, $\mathbf{t}^{*}$, $l^{*}$
+Appendix B Experiment Details
 B.1 Training Data
 The datasets used in our training are listed in Table 5. The number of frames may slightly differ from that of the original data because some invalid frames are dropped.
+To assign balanced weights to the datasets for training, we compute the retrieval probability of each dataset relative to OpenImagesV7 [32], a large and diverse natural image dataset.
+Specifically, we leverage DINOv2 [41] to extract feature vectors and calculate the probability that the nearest neighbor of a randomly selected image from OpenImagesV7 is found in each respective training dataset.
 Name Domain #Frames Type Weight
-Training
 A2D2[21] Outdoor/Driving $196$K C 0.8%
@@
 ObjaverseV1[12] Object $167$K A 4.8%
-Table 6: A table of datasets used for training.
-#Frames stands for the number of used frames.
-The types are in short for A.
-synthetic, B.
-structure from motion / multiview reconstruction, C.
-LiDAR / Laser; D.
-Microsoft Kinect
+Type Label quality Applied losses
+Accuracy Range Density $\mathcal{L}_{G}$ $\mathcal{L}_{S_{1}}$ $\mathcal{L}_{S_{2}}$ $\mathcal{L}_{S_{3}}$ $\mathcal{L}_{N}$ $\mathcal{L}_{M}$
+A. Synthetic Perfect $\infty$ Dense
+B. SfM/MV Recon High $\infty$ Dense&Partial
+C. LiDAR/Laser High $\sim 100$m Sparse
+D. Kinect Medium $\sim 10$m Dense
+Table 5: Datasets used for training (top) and the tailored loss combinations (bottom).
 B.2 Evaluation Data
 The raw evaluation datasets are processed accordingly for reliable evaluation and fair comparison. We report the details as follows.
-  • NYUv2 [38] We use the official test split of 654 samples Due to the inaccuracy of ground truth values captured by Kinect V1 near boundaries, we filter and remove boundary regions by a simple edge detection method.
+Figure 8: Examples of evaluation data preprocessing: (a) Removing mirror and boundary artifacts from the ground truth depth in NYUv2.
+(b) Excluding sky regions in Sintel.
+(c) Removing boundary artifacts from the ground truth depth in DIODE.
+  • NYUv2 [39].
+We use the official test split of 654 samples.
+Due to the inaccuracy of ground truth values captured by Kinect V1 near boundaries, we filter and remove boundary regions by a simple edge detection method.
 Specifically, we set a threshold for the difference between the minimum and maximum depth values within a local window.
-Depth values beyond 5 meters are excluded because they are unreliable due to the limited sensor range.
+Depth values beyond 5 meters are excluded because they are unreliable due to the limited sensor range [52].
 Additionally, we manually mask out areas with reflective and transparent surfaces, such as mirrors and glass, which cannot be accurately captured by the sensor.
-  • KITTI [51] We utilizes the test split of 652 images of Eigen et.
-al.
+  • KITTI [53].
+We utilize the test split of 652 images of Eigen et al.
 [18] following previous works.
-The original resolution of $1242\times 375$ does not match our training aspect ratio (ranging from $1:2$ to $2:1$), so we apply a center crop to obtain a $750\times 375$ resolution from the raw images.
-  • ETH3D [49] All 454 images are included. The images are undistorted with the official calibration data and downsized from the original $6202\times 4135$ resolution to $2048\times 1365$.
-  • iBims-1 [31] All 100 images are included at an original resolution of $640\times 480$.
-  • GSO [16] The dataset contains 1,030 objects.
+The original resolution of $1242\times 375$ does not match our training aspect ratio (ranging from $1:2$ to $2:1$), so we apply center cropping to obtain a resolution of $750\times 375$ from the raw images.
+  • ETH3D [50]. All 454 images are included. The images are undistorted with the official calibration data and downsized from the original resolution of $6202\times 4135$ to $2048\times 1365$.
+  • iBims-1 [31]. All 100 images are included at an original resolution of $640\times 480$.
+  • GSO [16].
+The dataset contains 1,030 objects.
 For each object, we render a single view at $512\times 512$ resolution.
@@
 The object is centered in the image, and its bounding box occupies approximately 70% of the image’s size.
-  • Sintel [9] We use all 1,064 frames and perform a center crop to $872\times 436$ from the original $1024\times 436$ resolution to fit our aspect ratio range.
-The sky regions are manually masked out.
-  • DDAD [24] We randomly select 1,000 samples from the validation set.
+  • Sintel [9].
+We use all 1,064 frames and center-crop the images to $872\times 436$ from the original $1024\times 436$ resolution to fit our aspect ratio range.
+The sky regions are manually masked out because evaluating models with sky depth included is not meaningful.
+  • DDAD [24].
+We randomly select 1,000 samples from the validation set.
 The dataset was collected using multiple cameras and LiDAR sensors mounted on a moving vehicle.
 Some cameras inadvertently capture parts of the vehicle, causing discrepancies with the sensor’s depth data.
-To address this, we crop the regions that are not obstructed by the vehicle itself.
-  • DIODE [52] We utilize the official validation split, which includes 325 indoor images and 446 outdoor images at an original resolution of $1024\times 768$.
+To address this issue, we crop the regions that are not obstructed by the vehicle itself.
+  • DIODE [54].
+We utilize the official validation split, which includes 325 indoor images and 446 outdoor images at an original resolution of $1024\times 768$.
 Due to artifacts in ground truth depth values near the boundaries in this dataset, we identify and remove these boundary regions using a similar approach as described above.
+B.3 Evaluation Protocol
+For all our models and baselines, predictions and ground truth are aligned in scale (and shift, if applicable) for each image before measuring errors. To clarify the notations in this section:
+  • $\hat{\mathbf{p}}_{i}$ and $\mathbf{p}_{i}$ are the predicted and ground-truth points, respectively.
+  • $\hat{z}_{i}$ and $z_{i}$ are the predicted and ground-truth depths, which are the $Z$-coordinate of corresponding points.
+  • $\mathcal{M}$ is the mask of valid ground-truth.
+  • $a$ and $b$ denote the scale and shift used to align predictions with the ground truth for evaluation, to avoid confusion with similar symbols used in the training objectives.
+  • Scale-invariant point map. The scale $a^{*}$ to align prediction with ground truth is computed as:
+    $a^{*}=\mathop{\arg\!\min}_{a}\sum_{i\in\mathcal{M}}{1\over z_{i}}\|a\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|_{1},$ (23)
+  • Affine-invariant point map. The scale $a^{*}$ and shift $\mathbf{b}^{*}$ are computed as:
+    $(a^{*},\mathbf{b}^{*})=\mathop{\arg\!\min}_{a,\mathbf{b}}\sum_{i\in\mathcal{M}}{1\over z_{i}}\|a\hat{\mathbf{p}}_{i}+\mathbf{b}-\mathbf{p}_{i}\|_{1}.$ (24)
+  • Scale-invariant depth map. The scale $a^{*}$ is computed as
+    $a^{*}=\mathop{\arg\!\min}_{s}\sum_{i\in\mathcal{M}}{1\over z_{i}}|a\hat{z}_{i}-z_{i}|.$ (25)
+  • Affine-invariant depth map. The scale $a^{*}$ and shift $b^{*}$ are computed as
+    $(a^{*},b^{*})=\mathop{\arg\!\min}_{s}\sum_{i\in\mathcal{M}}{1\over z_{i}}|a\hat{z}_{i}+b-z_{i}|.$ (26)
+  • Affine-invariant disparity map. We follow the established protocol for affine disparity alignment [43], using least-squares to align predictions in disparity space:
+    $(a^{*},b^{*})=\mathop{\arg\!\min}_{s}\sum_{i\in\mathcal{M}}(a\hat{d}_{i}+b-d_{i})^{2},$ (27)
+    where $\hat{d}_{i}$ is the predicted disparity and $d_{i}$ is the ground truth, defined as $d_{i}=1/z_{i}$.
+To prevent aligned disparities from taking excessively small or negative values, the aligned disparity is truncated by the inverted maximum depth $1/z_{\text{max}}$ before inversion.
+The final aligned depth $\hat{z}_{i}^{*}$ is computed as:
+    $\hat{z}_{i}^{*}:={1\over\max(a^{*}\hat{d}_{i}+b^{*},1/z_{\text{max}})}.$ (28)
 Appendix C More Results
-C.1 Full table of depth estimation results
-In Table 7, we present a full table of results for depth estimation where methods that predict metric or scale-invariant depth are evaluated for affine-invariant depth and disparity for fair comparison with other baselines.
+Full table of depth estimation results.
+In Table 6, we present detailed results for depth estimation where methods that predict metric or scale-invariant depth are also evaluated on affine-invariant depth and disparity for a fair comparison.
 Method NYUv2 KITTI ETH3D iBims-1 GSO Sintel DDAD DIODE Average
@@
 DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 3.06
-Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.944 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.38
+Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.38
 Table 6: Full table of comparison for depth map estimation.
-* Methods have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results in smaller text size for each version.
+*: methods that have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results for each version (smaller font size).
 Gray numbers denote models trained on respective benchmarks.
-C.2 Qualitative comparison on uncurated in-the-wild images
-We further utilize the first 100 in-the-wild images in DIV2K[2] dataset to extensively compare our method with LeReS [69], UniDepth [41], DUSt3R [57], Metric3D V2 [26] and Depth Anything V2 [66].
-The results can be found in the supplementary videos.
+More qualitative comparisons.
+Fig.
+10 and Fig.
+11 present additional visual comparisons on zero-shot evaluation datasets and in-the-wild images.
+Our method is compared with LeReS [71], UniDepth [42], DUSt3R [59], Metric3D V2 [26] and Depth Anything V2 [68].
 Since Metric3D V2 and Depth Anything V2 predict depth map and require ground truth camera focal to obtain 3D points cloud results, we visualize them using our estimated focal lengths.
-C.3 More visual results
-In the following pages, we demonstrate more examples of our results for in-the-wild images.
-The rows from left to right are the input images, the disparity maps, surface geometry visualization from the source view and three novel view image of the point cloud, respectively.
+In the supplementary videos, we present extensive and uncurated comparisons using the first 100 images from the DIV2K[2] dataset.
+More visual results
+In Fig. 12 and Fig. 13, we demonstrate more reconstruction results of our method for more open-domain images.
+Appendix D Limitations and Future Work
+While our model demonstrates strong performance, accurately capturing thin structures remains a significant challenge.
+This difficulty arises from the network’s limited capacity and the presence of noisy real-world training data.
+As illustrated in Fig.
+9, our model may fail to recover these intricate structures.
+Additionally, while monocular video reconstruction holds great promise as an application, achieving temporal consistency presents substantial challenges.
+Our model, designed for single-image input, cannot inherently maintain temporal coherence due to the ambiguity of the task.
+Addressing this issue would require non-trivial solutions, such as global optimization techniques.
+Given the rapid advancements in video depth estimation, we believe that an end-to-end model for monocular video reconstruction could significantly benefit from our proposed techniques.
+Exploring this direction is a compelling avenue for future work.
+Figure 9: A failure case. Our model does not capture the thin structures of the fence, leading to a flattened geometry.
+Figure 10: Additional qualitative comparisons from the evaluation datasets.
+*: for methods without camera intrinsics prediction, ground-truth camera intrinsics (and disparity shifts) were used to lift their results into 3D points.
+Best viewed with zoom.
+Figure 11: Additional qualitative comparisons for in-the-wild images from the DIV2K dataset.
+*: for methods without camera intrinsics prediction, our camera intrinsics prediction were used to lift their results into 3D points.
+Supplementary videos contain more results.
+Best viewed with zoom.
+Figure 12: Additional visual results for open-domain images of our model (page 1 of 2).
+The columns from left to right are the input images, reconstructed disparity maps, reconstructed surface geometry viewed from the source view, and three novel-view images, respectively.
+Figure 13: Additional visual results for open-domain images of our model (page 2 of 2).
+The columns from left to right are the input images, reconstructed disparity maps, reconstructed surface geometry viewed from the source view, and three novel-view images, respectively.

### v3

@@
 Scale-invariant point map
-LeReS 17.0 76.0 37.3 11.1 17.1 75.8 18.5 72.1 14.7 75.9 38.6 30.6 31.9 39.6 27.5 46.7 25.3 53.5 3.94
-DUSt3R 5.56 97.1 21.9 63.6 10.7 90.5 6.17 95.4 4.55 99.3 34.8 50.4 21.3 70.3 12.3 87.0 14.7 81.7 2.69
-UniDepth 5.36 98.3 10.7 98.0 18.5 77.5 5.27 97.4 6.57 99.6 33.0 49.0 11.3 90.3 12.3 91.1 12.9 87.7 2.18
-Ours 4.91 98.4 8.62 96.3 4.60 98.9 4.62 97.1 2.60 100 22.3 69.4 12.3 90.4 6.50 94.7 8.32 93.1 1.12
+LeReS 16.9 76.0 31.6 28.4 17.1 75.8 18.5 72.2 14.7 76.0 38.6 30.6 32.0 39.4 27.6 46.4 24.6 55.6 3.94
+DUSt3R 5.53 97.1 15.2 87.9 10.7 90.6 6.18 95.4 4.54 99.3 34.8 50.3 21.4 70.1 12.4 86.7 13.8 84.7 2.75
+UniDepth 5.33 98.4 5.96 98.5 18.5 77.6 5.29 97.4 6.58 99.6 33.0 48.9 11.4 90.2 12.3 91.0 12.3 87.7 2.09
+Ours 4.86 98.4 5.47 97.4 4.58 98.9 4.63 97.1 2.58 100 22.3 69.5 12.3 90.3 6.58 94.5 7.91 93.3 1.22
 Affine-invariant point map
-LeReS 9.55 91.4 29.0 43.8 14.7 79.6 11.0 88.7 8.89 95.3 29.7 55.5 29.3 46.8 14.9 80.3 18.4 72.7 3.94
-DUSt3R 4.49 97.4 18.0 66.7 7.27 94.9 5.02 96.0 3.08 99.5 30.3 56.7 19.6 71.4 8.86 88.9 12.1 83.9 2.94
-UniDepth 3.96 98.4 7.94 97.3 12.2 89.6 4.62 98.0 2.98 99.8 28.5 58.5 10.2 90.6 8.46 91.1 9.86 90.4 1.88
-Ours 3.73 98.3 7.69 95.0 3.60 99.0 3.60 97.3 1.16 100 16.9 77.8 10.4 91.5 4.31 96.6 6.43 94.4 1.25
+LeReS 9.51 91.4 26.1 49.1 14.7 79.6 11.0 88.6 8.91 95.2 29.7 55.5 29.4 46.7 15.1 80.1 18.1 73.3 3.94
+DUSt3R 4.45 97.4 12.7 83.3 7.27 95.0 5.04 96.0 3.07 99.6 30.3 56.6 19.7 71.2 8.97 88.7 11.4 86.0 2.94
+UniDepth 3.93 98.4 4.29 98.6 12.2 89.6 4.65 98.0 2.99 99.8 28.5 58.4 10.3 90.5 8.56 90.9 9.43 90.5 1.81
+Ours 3.68 98.3 4.86 97.2 3.57 99.0 3.61 97.3 1.14 100 16.8 77.8 10.5 91.4 4.37 96.4 6.07 94.7 1.31
 Local point map
-LeReS - - 8.62 91.5 15.7 84.4 - 18.7 85.8 1.65 99.0 16.2 85.2 12.2 89.2 3.70
-DUSt3R - - 5.88 94.4 9.43 91.3 - 12.7 88.3 1.71 98.9 10.1 91.3 7.96 92.9 2.30
-UniDepth - - 7.45 92.7 10.2 90.2 - 17.0 86.9 1.64 98.8 13.2 87.8 9.89 91.3 3.00
-Ours - - 2.86 97.9 7.02 94.1 - 9.73 92.1 1.08 99.5 5.96 95.2 5.33 95.8 1.00
+LeReS - - - - 9.32 91.9 8.57 93.2 - - 13.3 84.8 10.7 88.9 11.6 88.2 10.7 89.4 3.80
+DUSt3R - - - - 6.05 94.8 5.44 95.9 - - 11.8 87.0 9.24 90.8 7.32 93.1 7.97 92.3 2.30
+UniDepth - - - - 8.61 92.6 5.92 96.0 - - 13.4 84.3 8.18 92.0 9.95 90.0 9.21 91.0 2.90
+Ours - - - - 3.21 98.1 4.16 96.8 - - 8.63 92.7 6.74 94.3 4.78 96.3 5.50 95.6 1.00
 […]
@@
 Scale-invariant depth map
-LeReS 12.1 82.5 19.3 64.3 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.1 18.1 69.9 18.5 69.5 7.31
-ZoeDepth 5.65 96.3 7.23 92.2 10.4 87.3 7.48 93.2 3.23 99.9 27.4 61.7 17.0 73.0 11.2 85.4 11.2 86.1 5.50
-DUSt3R 4.43 97.1 7.71 90.9 6.07 95.6 4.96 95.9 3.27 99.5 31.1 57.2 18.5 73.5 8.84 88.8 10.6 87.3 5.00
-Metric3D V2 4.74 97.3 3.96 98.6 3.92 98.5 4.20 97.7 2.47 99.9 20.7 70.0 7.36 94.6 3.27 98.5 6.32 94.4 2.07
-UniDepth 3.88 98.3 3.73 98.7 5.68 97.0 4.76 97.4 4.17 99.7 28.3 58.9 9.98 90.7 6.79 92.9 8.41 91.7 3.00
-DA V1 4.81 97.5 5.52 95.8 8.80 92.3 5.54 95.8 5.48 99.2 28.1 56.7 13.1 81.7 10.1 87.8 10.2 88.4 5.67
-DA V2 5.09 97.2 7.28 93.6 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 14.6 78.2 7.89 90.3 9.08 89.7 4.06
-Ours 3.48 98.4 4.27 97.8 3.38 98.9 3.45 97.1 1.49 100 19.4 73.4 9.11 90.6 4.85 94.8 6.17 93.9 1.62
+LeReS 12.1 82.6 19.2 64.8 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.0 18.2 69.6 18.5 69.5 7.31
+ZoeDepth 5.62 96.3 7.27 91.9 10.4 87.3 7.45 93.2 3.23 99.9 27.4 61.8 17.0 72.8 11.3 85.2 11.2 86.1 5.50
+DUSt3R 4.40 97.1 7.81 90.6 6.04 95.7 4.98 95.8 3.27 99.5 31.1 57.2 18.6 73.3 8.91 88.8 10.6 87.2 5.00
+Metric3D V2 4.69 97.4 4.00 98.5 3.84 98.5 4.23 97.7 2.46 99.9 20.7 69.8 7.41 94.6 3.29 98.4 6.33 94.3 2.07
+UniDepth 3.86 98.4 3.73 98.6 5.67 97.0 4.79 97.4 4.18 99.7 28.3 58.8 10.1 90.5 6.83 92.8 8.43 91.6 3.00
+DA V1 4.77 97.5 5.61 95.6 9.41 88.9 5.53 95.8 5.49 99.3 28.3 56.7 13.2 81.5 10.3 87.5 10.3 87.9 5.67
+DA V2 5.03 97.3 7.23 93.7 6.12 95.5 4.32 97.9 4.38 99.3 23.0 65.2 14.7 78.0 7.95 90.0 9.09 89.6 4.06
+Ours 3.44 98.4 4.25 97.8 3.36 98.9 3.46 97.0 1.47 100 19.3 73.4 9.17 90.5 4.89 94.7 6.17 93.8 1.62
 Affine-invariant depth map
@@
 GeoWizard 4.69 97.4 8.14 92.5 6.90 94.0 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.44 90.7 2.69
-Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.00
+Ours 2.92 98.6 3.94 98.0 2.69 99.2 2.74 97.9 0.94 100 13.0 83.2 8.40 92.1 3.16 97.5 4.72 95.8 1.00
 Affine-invariant disparity map
-MiDaS V3.1 4.62 98.1 6.30 94.8 5.80 96.8 4.75 97.4 1.88 100 21.4 73.0 14.4 82.7 6.06 94.9 8.15 92.2 3.69
-DA V1 4.24 98.4 5.40 97.0 4.70 98.2 4.20 97.6 1.54 100 20.4 77.5 12.6 87.0 5.72 95.7 7.36 93.9 2.31
-DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 2.50
-Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.06
+MiDaS V3.1 4.58 98.1 6.25 94.7 5.77 96.8 4.73 97.4 1.86 100 21.3 73.1 14.5 82.6 6.05 94.9 8.13 92.2 3.69
+DA V1 4.20 98.4 5.40 97.0 4.68 98.2 4.18 97.6 1.54 100 20.1 77.6 12.7 86.9 5.69 95.7 7.31 93.9 2.31
+DA V2 4.14 98.3 5.61 96.7 4.71 97.9 3.47 98.5 1.24 100 21.4 72.8 13.1 86.4 5.29 96.1 7.37 93.3 2.56
+Ours 3.38 98.6 4.05 98.1 3.11 98.9 3.23 98.0 0.96 100 18.4 79.5 8.99 91.5 3.98 97.2 5.76 95.2 1.06
 Table 2: Quantitative results for depth map estimation.
 Gray numbers denote models trained on respective benchmarks and thus excluded from ranking.
-An extensive comparison is found in the suppl.
+A more extensive comparison can be found in the suppl.
 materials.
@@
 […]
-Depth+camera 11.3 88.7 9.20 90.4 10.1 91.2 8.97 90.1 7.29 92.6 8.22 92.1
-Opt. scale-inv. 10.5 89.7 8.50 91.4 9.09 92.0 8.28 90.9 6.72 93.3 7.91 92.6
-Opt. affine-inv. 10.1 90.3 8.07 91.9 7.84 93.2 7.92 91.2 6.29 93.7 7.43 93.2
-Med. affine-inv. 11.5 88.5 9.36 89.7 10.2 90.9 9.11 89.7 7.38 92.4 8.53 92.0
-Full w/o truncation 10.1 90.4 8.13 91.4 6.94 94.0 7.94 91.4 6.31 93.5 7.45 93.1
-Full w/o $\mathcal{L}_{S}$ 10.1 90.2 8.10 91.8 7.51 93.5 7.94 91.2 6.31 93.6 7.47 93.2
-Full 10.0 90.5 8.00 91.8 7.07 94.0 7.84 91.4 6.21 93.8 7.31 93.3
-Table 4: Quantitative ablation study results. All experiments are conducted with a ViT-Base backbone for the encoder.
+SI-Log depth 11.2 88.7 9.09 90.6 9.19 91.2 8.94 90.1 7.27 92.6 8.23 92.1
+Affine-inv. depth 29.9 51.4 29.0 52.7 12.2 86.0 28.9 52.7 6.18 93.9 15.9 76.6
+ROE scale-inv. 10.3 89.8 8.34 91.6 8.59 91.9 8.27 90.9 6.73 93.2 7.90 92.6
+L2 affine-inv. 13.5 84.2 10.3 88.2 9.48 91.0 11.1 85.7 8.03 91.2 9.37 90.5
+Med. affine-inv. 10.9 89.0 8.97 90.7 9.44 90.7 9.10 89.8 7.50 92.4 8.74 91.8
+ROE affine-inv. 9.84 90.3 7.88 92.1 7.62 93.3 7.91 91.2 6.29 93.7 7.43 93.2
+Full w/o trunc. 9.81 90.5 7.91 91.7 7.12 93.8 7.92 91.3 6.31 93.5 7.45 93.1
+Full w/o $\mathcal{L}_{S}$ 9.98 90.3 7.94 92.1 7.47 93.4 7.94 91.2 6.30 93.6 7.47 93.2
+Full 9.78 90.6 7.83 92.1 7.16 93.8 7.82 91.3 6.20 93.7 7.30 93.3
+Table 4: Quantitative ablation study results. All experiments are conducted with a ViT-Base encoder. The first six rows are trained with global loss only.
 Figure 6: Qualitative ablation study results.
@@
 The principal point of the cropped image is aligned with its center using a perspective crop.
-4.1 Quantitative Evaluations
+4.1 Quantitative Evaluations.
 We assess the zero-shot performance of our model and compare it to several state-of-the-art methods for monocular point map estimation [71, 42, 59] and depth map estimation [67, 68, 6, 26, 5, 20, 29].
@@
 As detailed below, our method achieves superior results across various tasks and benchmarks.
+[^1] In this revision, we corrected evaluation numbers due to a minor error in the evaluation data.
+The overall rankings and conclusions are not affected.
+All results are reproducible using the released code.
 Point map estimation.
@@
 material for more preprocessing details).
-For evaluation metrics, we use the relative point error Relp, i.e., $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}<0.25$.
+For evaluation metrics, we use the relative point error Relp, i.e., $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\min(\|\mathbf{p}\|_{2},\|\hat{\mathbf{p}}\|_{2})<0.25$.
 We compare our method with LeReS [71], UniDepth [42], and DUSt3R [59], evaluating both camera-space scale-invariant point maps and affine-invariant point maps.
@@
 Due to the need for high-quality ground truth and scenes with multiple objects, NYUv2, KITTI, and GSO are excluded.
-Our method achieves a substantial reduction in region-wise Relp, lowering the error from $7.96$ to $5.33$—an approximately $30\%$ improvement over previous methods.
+Our method achieves a substantial reduction in region-wise Relp, lowering the error from $7.97$ to $5.50$—an approximately $30\%$ improvement over previous methods.
 Depth map estimation.
@@
 Point map representation.
-We compare our affine-invariant point map representation with two variants, one scale-invariant point map in camera space, and another that predicts a depth map and camera parameters.
-For the former, we apply the proposed ROE alignment to solve the scale parameter only.
-For the latter, we adopt the SI-log loss [18] for the depth map and the MSE loss for the ray direction map, similar to UniDepth [42].
-Results are shown in the first three rows of Table 4, where only the global ground-truth supervision is applied in these experiments.
-Our proposed affine-invariant representation outperforms the others across various tasks.
-This is attributed to the fact that the other two variants suffer from focal-distance ambiguity issue during training, which hinders their performance.
+We compare our affine-invariant point map representation with three variants: (1) a scale-invariant depth map and a dense ray map for camera parameters, (2) an affine-invariant depth map with a shift scalar and a dense ray map, and (3) a scale-invariant point map in camera space.
+The first variant uses SI-log loss [18] for the depth map and MSE for the ray map, follows UniDepth [42].
+The second employs ROE alignment with $L_{1}$ loss for the affine depth, $L_{1}$ loss for shift, and MSE for the ray map.
+The third applies the ROE alignment to solve the scale parameter only.
+Table 4 (row 1-3) shows that our affine-invariant representation outperforms the others across various tasks.
+Notably, while the affine-invariant depth representation slightly improves specific metrics, all others drop dramatically.
+Our representation excels by effectively resolving focal-distance ambiguity, a key issue that hampers effectively training of the other three variants.
 ROE alignment.
-To evaluate the effectiveness of our optimal alignment, we compare it to a previously-used simple alignment method [43], i.e., the predicted and ground-truth point maps are normalized by their median: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})$ equals to $(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})$ is computed as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$.
-As shown in the third and fourth rows of Table 4, our proposed alignment method consistently outperforms the median alignment strategy across all tasks by large margins, demonstrating the critical role of our proposed optimal alignment.
-We ablate the truncation of residuals in alignment objects, which degrades performance (see the fifth row in Table 4). In practice, adding truncation also improves training stability.
+To assess the effectiveness of our optimal alignment, we compare it to least-squares alignment for $L_{2}$ loss and a previously-used median-based alignment [43] for $L_{1}$ loss.
+The former has an analytic solution, while the latter normalizes predicted and ground-truth point maps by their median: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})=(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})=\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$.
+As shown in the Table 4 (row 4-6), our proposed alignment method consistently outperforms the other two strategies across all tasks by large margins, highlighting its critical role.
+We also ablate the truncation of alignment objective residuals, which degrades performance (Table 4 row 7).
 Multi-scale local geometry loss.
-As shown in the six-th row of Table 4, the quantitative results without $\mathcal{L}_{S}$ suffer a significant drop on local geometry metrics.
+As shown in Table 4 (row 8), the quantitative results without $\mathcal{L}_{S}$ suffer a significant drop on local geometry metrics.
 This demonstrates the effectiveness of $\mathcal{L}_{S}$ in enhancing local geometry accuracy and compensating for the insufficient supervision of the global loss due to the ambiguous relative positions between distant scene objects.
@@
 D. Kinect Medium $\sim 10$m Dense
-Table 5: Datasets used for training (top) and the tailored loss combinations (bottom).
+Table 5: Datasets used for training and tailored loss combination.
 B.2 Evaluation Data
@@
     $(a^{*},\mathbf{b}^{*})=\mathop{\arg\!\min}_{a,\mathbf{b}}\sum_{i\in\mathcal{M}}{1\over z_{i}}\|a\hat{\mathbf{p}}_{i}+\mathbf{b}-\mathbf{p}_{i}\|_{1}.$ (24)
-  • Scale-invariant depth map. The scale $a^{*}$ is computed as
+  • Scale-invariant depth map, the scale $a^{*}$ is computed as
     $a^{*}=\mathop{\arg\!\min}_{s}\sum_{i\in\mathcal{M}}{1\over z_{i}}|a\hat{z}_{i}-z_{i}|.$ (25)
@@
 Appendix C More Results
-Full table of depth estimation results.
+Full table of depth estimation results
 In Table 6, we present detailed results for depth estimation where methods that predict metric or scale-invariant depth are also evaluated on affine-invariant depth and disparity for a fair comparison.
@@
 Scale-invariant depth
-LeReS 12.1 82.5 19.3 64.3 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.1 18.1 69.9 18.5 69.5 7.31
-ZoeDepth 5.65 96.3 7.23 92.2 10.4 87.3 7.48 93.2 3.23 99.9 27.4 61.7 17.0 73.0 11.2 85.4 11.2 86.1 5.50
-DUSt3R 4.43 97.1 7.71 90.9 6.07 95.6 4.96 95.9 3.27 99.5 31.1 57.2 18.5 73.5 8.84 88.8 10.6 87.3 5.00
-Metric3D V2 4.74 97.3 3.96 98.6 3.92 98.5 4.20 97.7 2.47 99.9 20.7 70.0 7.36 94.6 3.27 98.5 6.32 94.4 2.07
-UniDepth 3.88 98.3 3.73 98.7 5.68 97.0 4.76 97.4 4.17 99.7 28.3 58.9 9.98 90.7 6.79 92.9 8.41 91.7 3.00
-DA V1* 4.81 97.5 5.52 95.8 8.80 92.3 5.54 95.8 5.48 99.2 28.1 56.7 13.1 81.7 10.1 87.8 10.2 88.4 5.67
- - metric indoor 4.81 97.5 15.2 74.2 9.42 89.0 5.54 95.8 5.48 99.2 28.3 56.7 24.2 57.5 10.1 87.8 12.9 82.2 -
- - metric outdoor 16.0 72.2 5.52 95.8 8.80 92.3 13.8 78.8 8.60 93.6 28.1 54.8 13.1 81.7 12.9 81.7 13.3 81.4 -
-DA V2* 5.09 97.2 7.28 93.6 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 14.6 78.2 7.89 90.3 9.08 89.7 4.06
- - metric indoor 5.09 97.2 7.49 91.3 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 16.5 73.6 7.89 90.3 9.35 88.8 -
- - metric outdoor 15.3 72.1 7.28 93.6 9.33 89.6 10.6 84.9 9.61 92.5 28.6 57.4 14.6 78.2 12.1 83.6 13.4 81.5 -
-Ours 3.48 98.4 4.27 97.8 3.38 98.9 3.45 97.1 1.49 100 19.4 73.4 9.11 90.6 4.85 94.8 6.17 93.9 1.62
+LeReS 12.1 82.6 19.2 64.8 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.0 18.2 69.6 18.5 69.5 7.31
+ZoeDepth 5.62 96.3 7.27 91.9 10.4 87.3 7.45 93.2 3.23 99.9 27.4 61.8 17.0 72.8 11.3 85.2 11.2 86.1 5.50
+DUSt3R 4.40 97.1 7.81 90.6 6.04 95.7 4.98 95.8 3.27 99.5 31.1 57.2 18.6 73.3 8.91 88.8 10.6 87.2 5.00
+Metric3D V2 4.69 97.4 4.00 98.5 3.84 98.5 4.23 97.7 2.46 99.9 20.7 69.8 7.41 94.6 3.29 98.4 6.33 94.3 2.07
+UniDepth 3.86 98.4 3.73 98.6 5.67 97.0 4.79 97.4 4.18 99.7 28.3 58.8 10.1 90.5 6.83 92.8 8.43 91.6 3.00
+DA V1 4.77 97.5 5.61 95.6 9.41 88.9 5.53 95.8 5.49 99.3 28.3 56.7 13.2 81.5 10.3 87.5 10.3 87.9 5.67
+  -metric indoor 4.77 97.5 15.4 73.6 9.41 88.9 5.53 95.8 5.49 99.3 28.3 56.7 24.2 57.4 10.3 87.5 12.9 82.1 -
+  -metric outdoor 15.9 72.3 5.61 95.6 8.77 92.4 13.8 78.8 8.59 93.6 28.1 54.8 13.2 81.5 13.0 81.4 13.4 81.3 -
+DA V2 5.03 97.3 7.23 93.7 6.12 95.5 4.32 97.9 4.38 99.3 23.0 65.2 14.7 78.0 7.95 90.0 9.09 89.6 4.06
+  -metric indoor 5.03 97.3 7.61 90.9 6.12 95.5 4.32 97.9 4.38 99.3 23.0 65.2 16.6 73.4 7.95 90.0 9.38 88.7 -
+  -metric outdoor 15.3 72.3 7.23 93.7 9.30 89.6 10.6 84.9 9.62 92.5 28.6 57.3 14.7 78.0 12.2 83.2 13.4 81.4 -
+Ours 3.44 98.4 4.25 97.8 3.36 98.9 3.46 97.0 1.47 100 19.3 73.4 9.17 90.5 4.89 94.7 6.17 93.8 1.62
 Affine-invariant depth
-LeReS 6.25 95.4 8.15 90.6 8.98 90.7 6.71 94.6 4.03 99.4 24.0 64.8 16.1 75.9 9.93 88.4 10.5 87.5 8.81
-ZoeDepth 4.79 97.2 5.51 95.3 7.29 94.2 5.88 95.7 2.54 99.9 21.8 69.2 14.1 80.3 7.72 91.0 8.70 90.3 7.25
-DUSt3R 3.76 97.8 7.21 91.8 4.96 96.4 3.92 96.6 2.56 99.6 25.4 64.1 16.8 76.4 6.55 92.9 8.89 89.5 6.69
-Metric3D V2 4.00 97.6 3.48 98.5 3.54 98.4 3.25 98.3 2.12 99.4 24.4 71.6 7.10 94.8 2.73 98.8 6.33 94.7 3.50
-UniDepth 3.42 98.6 3.54 98.8 4.93 97.5 3.74 98.2 2.47 99.9 24.8 64.1 9.36 91.0 4.88 96.2 7.14 93.0 3.62
-Marigold 4.63 97.3 7.29 93.8 6.08 96.3 4.35 97.2 2.78 99.9 21.2 75.0 14.6 80.5 6.34 94.3 8.41 91.8 5.81
-GeoWizard* 4.69 97.4 8.14 92.5 6.90 94.0 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.44 90.7 6.56
-- indoor 4.69 97.4 8.14 92.5 6.90 93.9 4.50 97.1 2.00 99.9 17.8 76.0 16.5 75.7 7.03 92.7 8.44 90.7 -
-- outdoor 4.83 97.3 8.40 92.2 7.18 93.7 4.59 97.0 2.03 99.9 17.8 76.2 17.1 74.4 7.22 92.5 8.64 90.4 -
-- object 4.70 97.4 8.21 92.4 6.92 94.0 4.52 97.0 2.05 99.9 17.9 76.0 16.8 75.1 7.13 92.6 8.52 90.5 -
-DA V1* 3.86 98.3 4.94 96.6 6.26 96.5 4.24 97.3 1.98 100 20.1 71.8 11.2 86.2 6.50 93.4 7.55 92.8 4.50
- - metric indoor 3.86 98.3 9.78 87.2 6.35 95.0 4.24 97.3 1.98 100 20.1 71.8 16.9 74.2 6.50 93.0 8.72 89.6 -
- - metric outdoor 7.72 93.8 4.94 96.6 6.26 96.5 7.01 94.2 2.78 99.8 20.7 70.0 11.2 86.2 6.93 93.4 8.44 91.3 -
-DA V2* 4.22 97.8 6.81 94.2 4.66 97.2 3.45 98.3 1.45 100 17.2 76.5 14.2 82.0 5.32 94.9 7.16 92.6 3.19
- - metric indoor 4.22 97.8 6.95 92.7 4.66 97.2 3.45 98.3 1.45 100 17.2 76.5 14.2 80.0 5.32 94.9 7.18 92.2 -
- - metric outdoor 8.69 90.9 6.81 94.2 7.26 93.5 6.81 93.5 2.28 100 22.5 67.0 13.2 82.0 8.00 91.2 9.45 89.0 -
-Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.56
+LeReS 6.21 95.4 8.28 90.3 8.95 90.8 6.68 94.5 4.03 99.4 24.0 64.8 16.2 75.8 9.99 88.1 10.5 87.4 8.81
+ZoeDepth 4.76 97.3 5.59 95.1 7.27 94.2 5.85 95.7 2.54 99.9 21.8 69.2 14.2 80.1 7.80 90.9 8.73 90.3 7.33
+DUSt3R 3.73 97.8 7.30 91.6 4.96 96.4 3.94 96.6 2.55 99.6 25.4 64.2 16.9 76.2 6.68 92.6 8.93 89.4 6.62
+Metric3D V2 3.94 97.6 3.50 98.4 3.24 99.0 3.28 98.3 2.10 99.4 26.6 71.7 7.15 94.8 2.75 98.7 6.57 94.7 3.64
+UniDepth V1 3.40 98.6 3.55 98.7 4.92 97.5 3.76 98.2 2.48 99.9 24.9 64.1 9.46 90.8 4.90 96.2 7.17 93.0 3.62
+Marigold 4.63 97.3 7.29 93.8 6.08 96.3 4.35 97.2 2.78 99.9 21.2 75.0 14.6 80.5 6.34 94.3 8.41 91.8 5.69
+Geowizard 4.69 97.4 8.14 92.5 6.90 93.9 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.45 90.7 6.44
+DA V1 3.82 98.3 5.04 96.4 6.23 95.2 4.23 97.3 1.98 100 20.1 71.8 11.3 86.1 6.75 92.6 7.43 92.2 4.83
+  -metric indoor 3.82 98.3 9.95 86.5 6.23 95.2 4.23 97.3 1.98 100 20.1 71.8 17.0 74.0 6.75 92.6 8.76 89.5 -
+  -metric outdoor 7.68 93.8 5.04 96.4 6.21 96.6 7.00 94.2 2.77 99.8 20.6 70.0 11.3 86.1 7.03 93.2 8.45 91.3 -
+DA V2 4.16 97.9 6.77 94.3 4.63 97.2 3.44 98.3 1.44 100 17.1 76.6 13.4 81.8 5.41 94.6 7.04 92.6 -
+  -metric indoor 4.16 97.9 7.09 92.3 4.63 97.2 3.44 98.3 1.44 100 17.1 76.6 14.3 79.8 5.41 94.6 7.20 92.1 -
+  -metric outdoor 8.65 91.0 6.77 94.3 7.24 93.5 6.80 93.5 2.29 100 22.4 67.1 13.4 81.8 8.19 90.7 9.47 89.0 2.94
+Ours 2.92 98.6 3.94 98.0 2.69 99.2 2.74 97.9 0.94 100 13.0 83.2 8.40 92.1 3.16 97.5 4.72 95.8 1.56
 Affine-invariant disparity
-LeReS 7.33 95.4 12.4 86.7 10.2 90.1 8.45 92.9 4.33 99.7 28.9 59.6 23.3 73.1 10.8 88.3 13.2 85.7 8.25
-ZoeDepth 5.24 97.7 5.86 95.6 8.08 94.0 6.22 96.1 2.59 99.9 26.9 66.3 14.0 81.8 8.17 92.0 9.63 90.4 6.75
-DUSt3R 4.27 98.1 7.67 92.2 5.60 96.2 4.47 96.7 2.64 99.8 39.9 56.8 17.4 76.3 7.09 92.8 11.1 88.6 6.75
-Metric3D V2 13.4 81.4 3.69 98.3 4.38 97.6 8.56 92.3 1.81 100 21.8 72.5 7.32 94.2 7.62 90.4 8.57 90.8 5.29
-UniDepth 3.80 98.7 3.64 98.7 5.34 97.2 4.04 98.1 2.55 99.9 28.5 60.9 9.83 89.3 5.94 95.5 7.96 92.3 3.62
-MiDaS V3.1 4.62 98.1 6.30 94.8 5.80 96.8 4.75 97.4 1.88 100 21.4 73.0 14.4 82.7 6.06 94.9 8.15 92.2 5.00
-DA V1 4.24 98.4 5.40 97.0 4.70 98.2 4.20 97.6 1.54 100 20.4 77.5 12.6 87.0 5.72 95.7 7.36 93.9 3.00
-DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 3.06
-Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.38
+LeReS 7.31 95.5 12.2 87.1 10.2 90.1 8.44 92.9 4.33 99.7 28.9 59.6 23.4 73.0 10.7 88.3 13.2 85.8 8.25
+ZoeDepth 5.21 97.7 5.84 95.6 8.07 94.0 6.19 96.1 2.60 99.9 26.9 66.3 14.1 81.7 8.17 92.0 9.63 90.4 6.75
+DUSt3R 4.24 98.1 7.72 92.1 5.60 96.2 4.49 96.6 2.63 99.8 40.0 56.7 17.4 76.2 7.10 92.8 11.1 88.6 6.75
+Metric3D V2 13.4 81.5 3.76 98.2 4.30 97.7 8.55 92.3 1.80 100 21.8 72.4 7.35 94.1 7.70 90.2 8.58 90.8 5.29
+UniDepth V1 3.78 98.7 3.64 98.7 5.34 97.2 4.06 98.1 2.56 99.9 28.6 60.7 9.94 89.1 5.95 95.5 7.98 92.2 3.62
+MiDaS V3.1 4.58 98.1 6.25 94.7 5.77 96.8 4.73 97.4 1.86 100 21.3 73.1 14.5 82.6 6.05 94.9 8.13 92.2 5.00
+DA V1 4.20 98.4 5.40 97.0 4.68 98.2 4.18 97.6 1.54 100 20.1 77.6 12.7 86.9 5.69 95.7 7.31 93.9 3.00
+DA V2 4.14 98.3 5.61 96.7 4.71 97.9 3.47 98.5 1.24 100 21.4 72.8 13.1 86.4 5.29 96.1 7.37 93.3 3.12
+Ours 3.38 98.6 4.05 98.1 3.11 98.9 3.23 98.0 0.96 100 18.4 79.5 8.99 91.5 3.98 97.2 5.76 95.2 1.44
 Table 6: Full table of comparison for depth map estimation.
-*: methods that have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results for each version (smaller font size).
+* methods have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results in smaller text size for each version.
 Gray numbers denote models trained on respective benchmarks.
-More qualitative comparisons.
+More qualitative comparisons
 Fig.
@@
 Exploring this direction is a compelling avenue for future work.
-Figure 9: A failure case. Our model does not capture the thin structures of the fence, leading to a flattened geometry.
+Figure 9: A failure case. Our model fails to capture the thin structure of the fence, leading to a flattened geometry.
 Figure 10: Additional qualitative comparisons from the evaluation datasets.